### v3

# On the Robustness of Verbal Confidence of LLMs in Adversarial Attacks

Abstract
Robust verbal confidence generated by large language models (LLMs) is crucial for the deployment of LLMs to help ensure transparency, trust, and safety in many applications, including those involving human-AI interactions. In this paper, we present the first comprehensive study on the robustness of verbal confidence under adversarial attacks. We introduce attack frameworks targeting verbal confidence scores through both perturbation and jailbreak-based methods, and demonstrate that these attacks can significantly impair verbal confidence estimates and lead to frequent answer changes. We examine a variety of prompting strategies, model sizes, and application domains, revealing that current verbal confidence is vulnerable and that commonly used defence techniques are largely ineffective or counterproductive. Our findings underscore the need to design robust mechanisms for confidence expression in LLMs, as even subtle semantic-preserving modifications can lead to misleading confidence in responses.
1 Introduction
With the unprecedented advancement of large language models (LLMs) [llama3modelcard, minaee2024largelanguagemodelssurvey, OpenAI2024] and their widespread deployment in real-world systems, the robustness in acquiring LLMs’ confidence is crucial for a wide range of applications [chen-mueller-2024-quantifying, Geng2023ASO, xiong2024can, zhang2023pacelmpromptingaugmentationcalibrated] including high-stakes tasks [Omar2024.08.11.24311810, XU2025103455]. For human beings, the expression of epistemic uncertainty using natural language is an inherent ability. As the intelligence of AI systems advances rapidly, we envision in the future it will be important for AI systems to express their confidence verbally and accurately to produce seamless interactions in human-AI collaborations.
The ability of LLMs to verbally convey their confidence [geng2024survey], referred to as verbal confidence in this paper, is not only important for applications, but also an indicator of the intelligence of LLMs. Ensuring accurate and robust verbal confidence also underlies the objective of aligning AI models’ core values with those of humans, such as honesty [li2024survey, Yang2023AlignmentFH], which could be critical to avoiding mistrust [dhuliawala-etal-2023-diachronic, 10.1145/3491102.3501967, 10.1145/3351095.3372852] and maintaining future AI to be safe.
Figure 1: (a) Demonstration of how perturbations on prompts can affect final verbal confidence
(b) An overview of our proposed attack framework centered on using generated confidence scores to optimize attacks that either perturb the prompt or append a series of optimized trigger tokens that ultimately result in a reduction in confidence.
Many state-of-the-art (SOTA) LLMs do not allow direct access to model information which remains an essential limitation for confidence estimation [XU2025103455, zhang2023pacelmpromptingaugmentationcalibrated]. Even if they are available, logit-based confidences centered on attaining tokens probabilities and the aggregation of such scores may not reflect the true model confidence on the overall answer. Verbal confidence encourages AI systems to express their estimation of uncertainty and has attracted a considerable amount of recent research, which examine the mechanism and impact of LLMs’ verbal confidence [Geng2023ASO, groot-valdenegro-toro-2024-overconfidence, kumar-etal-2024-confidence, lin2022teaching, mielke-etal-2022-reducing, portillowightman2023strength, sun2024confidenceestimationllmbaseddialogue, DBLP:journals/corr/abs-2305-14975, xiong2024can, zhou-etal-2023-navigating], including the influence and implications on trustworthiness and transparency [huang2024trustllmtrustworthinesslargelanguage]. Numerous real-world industry systems have started to utilize verbal confidence [chen-mueller-2024-quantifying, macdonald2024generative, zhang2023pacelmpromptingaugmentationcalibrated], motivating the analysis of LLMs’ behaviour on this front as it is key for their successful operation.
Our work presents the first comprehensive study on the robustness of verbal confidence under adversarial attacks and addresses the basic question: how robust is the verbal confidence of existing large language models? We highlight the differences in how models perceive and manage uncertainty cues compared to how humans naturally understand them. Furthermore, given the ease of attaining verbal confidence from even the most guarded black-box models, it is key to determine how effectively these can be leveraged by adversaries as an attack objective to alter model behaviour in other ways, such as predicted answers.
To assess the Achilles’ heel of LLMs in expressing their confidence [carlini2023are, QIU2022278, REN2020346, wei2023jailbroken, zou2023universal], we propose a framework designed for attacking verbal confidence scores through both perturbation and jailbreak-based approaches, and show that these attacks can significantly jeopardize verbal confidence estimates. We explore three key questions for model deployment that have not been explored: (i) How vulnerable are these estimates under a range of adversarial attack techniques? (ii) How can these attacks be effectively crafted to target confidence and answer changes? (iii) How easily can such attacks be detected or neutralized with existing defence methods? Detailed discussions on real-world applications and the importance of maintaining proper confidence can be found in Appendix A.
In summary, the main contributions of our work are as follows:
  • To the best of our knowledge, this is the first work to provide a comprehensive study of adversarial attacks on verbal confidence (as shown in Figure 1) and demonstrate their threat to generated confidences and answers.
  • We design novel verbal confidence attacks based on both perturbation-based and jailbreak-based approaches, showing they can cause up to a 30% reduction in average confidence on user queries and induce a high rate of answer changes (up to 100% in originally correctly predicted samples).
  • We provide a thorough analysis of the stability and behaviour of verbal confidence in various scenarios and under different mitigation strategies, demonstrating that most mitigation strategies are ineffective or counterproductive against such confidence attacks.
2 Related Work
Adversarial Attacks in NLP. Adversarial attacks have served as a critical tool to reveal key vulnerabilities of vision and NLP models [10.1145/3593042, huq2020adversarial, QIU2022278]. Adversarial attack methods can be divided into perturbation-based and jailbreak-based attacks. Recent works of the former type have examined the robustness of LLMs under word-level attacks that target prediction and generation accuracy  [wang2023on, zhu2023promptbench] through threat vectors such as demonstrations [wang2023adversarial]. Jailbreak-based attacks involve manually or algorithmically [carlini2023are, wei2023jailbroken] generated trigger phrases that are appended to the input prompt to bypass safety guardrails [Deng2023MASTERKEYAJ]. Early methods utilized means such as optimized universal adversarial triggers (UAT) [wallace-etal-2019-universal] and top-$k$ gradient-based search [shin-etal-2020-autoprompt]. Recent variants of trigger attacks include the GCG attack [zou2023universal], black-box only approaches [lapid2023open], and those with different optimization methods such as genetic algorithms [liu2024autodan], policy gradient methods [diao2023blackbox], and using specially trained LLMs [paulus2024advprompterfastadaptiveadversarial].
Vulnerability of Model Confidence.
Compared to attacks that target victim models’ accuracy, adversarial attacks against models’ confidence has been understudied but have become more prominent as AI models are deployed into more applications where calibration robustness needs to investigated (see further discussion in Appendix A). Previous works primarily targeted logit-based confidence. For example, galil2021disrupting formulated confidence attacks and emde2023certified examined confidence attacks through certified robustness guarantees. obadinma2024calibration designed a framework that tunes adversarial attacks to target miscalibration. zeng2024uncertaintyfragilemanipulatinguncertainty examined attacks that target uncertainties derived from output tokens’ logits. None of these approaches, however, have investigated the vulnerability of verbal confidence. This, as discussed above, is a fundamental problem warranting further investigation.
Natural-Language Confidence.
Recently, a significant amount of research has been conducted to examine the effectiveness of explicitly generating confidence scores alongside natural language responses (see geng2024survey for a detailed survey). For example, lin2022teaching examined fine-tuning decoder-based LLMs to generate calibrated confidence. DBLP:journals/corr/abs-2305-14975 evaluated prompting models for confidence scores and gauge the calibration accordingly. xiong2024can examined multiple prompting, sampling, and aggregation strategies for verbal confidence scores. To the best of our knowledge, this research is the first to examine the robustness of verbal confidence through adversarial attacks.
3 Adversarial Attacks on Verbal Confidence
We investigate the robustness of verbal confidence produced by LLMs. Given a user prompt consisting of a query $\bm{\mathbf{X}}\equiv\{x_{1},\dots,x_{N}\}$ with a length of $N$ word tokens, users of an LLM can obtain a confidence level by providing the LLM with an additional task prompt $\bm{\mathcal{P}}\equiv\{p_{1},\dots,p_{M}\}$ of $M$ word tokens, which describes how the LLM should provide a confidence score and may give an illustrative demonstration of how to do it (i.e., through an example). The queried LLM then generates a sequence of tokens $\bm{\mathcal{Y}}\equiv\{y_{1},\ldots,y_{R}\}$ consisting of $R$ word tokens as a response, containing a subsequence representing its overall answer $\mathcal{\bm{A}}\subseteq\bm{\mathcal{Y}}$ to the query. A corresponding confidence score, which we call the verbal confidence $\mathcal{C}$, is a scalar quantity attained with the help of the task-specific prompt that reflects the model’s belief in the correctness of its answer $\mathcal{A}$. The score $\mathcal{C}$ is derived strictly from the output word tokens $\bm{\mathcal{Y}}$ alone without any knowledge of internal model states. We refer to the entire process as confidence elicitation (CE). Formally, confidence elicitation methods (CEMs) are functions that map a generated sequence of word tokens $\bm{\mathcal{Y}}$ to a confidence score, $\mbox{CEM}:\bm{\mathcal{Y}}\rightarrow\mathcal{C}$, given a specific LLM that serves as a mapping function from the input prompts to the output sequence $\mbox{LLM}:(\mathbf{X},\bm{\mathcal{P}})\rightarrow\bm{\mathcal{Y}}$. As detailed in Section 3.4, the specific CEM functions can be implemented with different prompting strategies, including Chain-of-Thought and Multi-Step prompting.
3.1 Verbal Confidence Attack Methods
We propose a series of verbal confidence attack (VCA) methods to automatically generate adversarial prompts. As Figure 1(a) demonstrates, a common range of alterations in a prompt can significantly vary confidence estimates or predicted answers. This is undesirable since confidence should be robust when the meaning of $\bm{\mathbf{X}}$ or $\bm{\mathcal{P}}$ is preserved. The framework for the proposed attacks can be seen in Figure 1(b), which presents our attack objectives based on generated numeric confidence scores.
[…]
3.2 Perturbation-based Confidence Attack
We propose two perturbation-based verbal confidence attacks: VCA-TF and VCA-TB. These VCA attacks are based on two popular black-box attack algorithms: Textfooler (TF) [DBLP:conf/aaai/JinJZS20] and Textbugger (TB) [DBLP:conf/ndss/LiJDLW19]. However, unlike the original algorithms, we design new scoring functions which are optimized using verbal confidence in contrast to class-wise probabilities derived from model logits. These methods first generate importance scores for each token in the sequence by determining the corresponding difference in the predicted confidence without the token present in the sequence. Then, each token is ranked and perturbations are prioritized on tokens with the highest difference. VCA-TF focuses on synonym substitutions, while VCA-TB also includes character modifications like common typos.
In addition, we formulate two attack types that iteratively attempt corrupting an input through random common typos (Typos/Ty.), or through word level modifications that include random synonym substitutions, adjacent token swaps, or token removals (SubSwapRemove/SSR) to optimize the attack objective. Detailed descriptions of these perturbation-based attacks can be found in Appendix B.
Algorithm 1 ConfidenceTriggers Optimization
1:Input: Set of confidence estimation methods $\mathbf{Z}$, Set of training prompts $\mathbf{X}$, # generations $G$, # prompts $\alpha$, # tokens per prompt $\beta$, # samples for initial estimation $S$, # training samples per iteration $\xi$, # elites $E$
2:Output: Trigger Token Sequence $\mathbf{T}$ of length $\beta$
3:for all $z\in\mathbf{Z}$ do
4:  Random subset sample $S$ examples from $\mathbf{X}$
5:  $\mathcal{L}_{z}$ $\leftarrow$ AverageConfidence($z$, $S$) $\triangleright$ w/o any trigger
6:end for
[…]
8:for $g=1$ to $G$ do
9:  for all $p\in$ $\bf{P}$ do $\triangleright$ Evaluate fitness of each prompt
10:   for all $z\in\mathbf{Z}$ do
11:     Random subset sample $\xi$ examples from $\mathbf{X}$
12:     $\mathcal{L}_{p_{g}}$ $\leftarrow\mathcal{L}_{p_{g}}\mathbin{\|}$ AverageConfidence($p$, $z$, $\xi$) - $\mathcal{L}_{z}$
13:   end for
14:   $\mathcal{L}_{p_{g}}\leftarrow$ $\sum\mathcal{L}_{p_{g}}$ / $|Z|$
15:  end for
16:  Sort $\mathbf{P}$ by $\mathcal{L}_{p_{g}}$
17:  $\mathbf{P}_{g}$ $\leftarrow$ [][\;]
18:  Add $E$ lowest loss prompts $p$ to $\mathbf{P}_{g}$
19:  while $|\mathbf{P}_{g}|$ << $|\mathbf{P}|$ do
[…]
21:                according to $\mathcal{L}_{p_{g}}$ Twice
22:   Choose random split point $v$
23:   $p_{ADV_{1}}$, $p_{ADV_{2}}$ $\leftarrow$ Cross-Over($p_{ADV_{1}}$$p_{ADV_{2}}$, $v$)
24:   idx $\leftarrow$ Randomly choose $20\%$ indices to mutate
25:   $p_{ADV_{3}}$ $\leftarrow$ Mutation($p_{ADV_{1}}$, $p_{ADV_{2}}$,idx)
26:   $\mathbf{P}_{g}$ $\leftarrow\mathbf{P}_{g}\mathbin{\|}$ $p_{ADV_{1}}$, $p_{ADV_{2}}$, $p_{ADV_{3}}$
27:  end while
28:  Keep first $\alpha$ prompts in $\mathbf{P}_{g}$
29:  $\mathbf{P}$ $\leftarrow$ $\mathbf{P}_{g}$
30:end for
31:for all $p\in$ $\bf{P}$ do
32:  for all $z\in\mathbf{Z}$ do
33:   Random subset sample $S$ cases from $\mathbf{X}$
34:   $\mathcal{L}_{p_{G}}$ $\leftarrow\mathcal{L}_{p_{G}}\mathbin{\|}$ AverageConfidence($p$, $z$, $S$)
35:  end for
36:  $\mathcal{L}_{p_{G}}\leftarrow$ $\sum\mathcal{L}_{p_{G}}$ / $|Z|$
37:end for
[…]
3.3 Jailbreak-based Confidence Attack
3.3.1 ConfidenceTriggers
In Figure 1(b), we show the potential for a more generalizable form of attack by designing a VCA method called ConfidenceTriggers, which is designed to target confidence based on a popular jailbreak approach [lapid2023open]. By optimizing a single series of trigger tokens that get appended to any number of prompts, LLMs are led to generate lowered confidences scores for these prompts. Our work is the first to optimize and study triggers that target verbal confidence scores.
Triggers are optimized using a completely black-box genetic algorithm (Algorithm 1). No model-specific information or tokenizers are required, corresponding to the setup that only needs verbal confidence. To optimize for confidence reduction, we first generate an initial set of $\alpha$ prompts consisting of $\beta$ number of random words. These words come from a list of approximately 2,000 words related to the topic of uncertainty to create an initial focused search space. This initialization process is accomplished using the InitializePrompts function in the algorithm. We append the triggers to the system prompt, as we will later show it is a more effective attack vector. This also has the byproduct that if a system or API is infiltrated with such a trigger, then confidence estimates on subsequent (benign) user queries are compromised. The loss function to be minimized is the difference in average verbal confidence (see AverageConfidence function in Algorithm 5) across $\xi$ randomly sampled training examples when using the tested trigger compared to the average confidence estimated on $S$ randomly sampled examples without the trigger.
To enable generalizability across different CEMs, we use four different variants to estimate the loss for each trigger (detailed below in Section 3.4). To produce prompts for the next generation, first $E$ number of elites are selected, then pairs of parent triggers (from the previous generation) are iteratively selected through top-$k$ tournament selection [10.1162/evco.1996.4.4.361], with $k=2$ to undergo Single Point Cross-Over [10.7551/mitpress/1090.001.0001] and Random (Uniform) Mutation [10.5555/534133] until there are $\alpha$ child prompts. For Cross-Over, a random split point is chosen and two child prompts are created whereby the first consists of the sequence from the first parent up to the split point, and then from the second parent after the split point, and vice versa for the second child. For Mutation, a third child is created by iteratively sampling token randomly from either the first parent or second parent. To be specific, 20% of indices are chosen as mutation points, where a random word is sampled instead. After $G$ generations, the lowest loss prompt is selected. The efficiency of this attack algorithm is dependent on the input parameters in Algorithm 1, with the equation for number of model calls needed to optimize a set of triggers being approximately $|\mathbf{Z}|*S+G*\xi*|\mathbf{Z}|*\alpha+\alpha*|\mathbf{Z}|*S$, although once optimized a set of tokens can be reused indefinitely. Detailed descriptions of settings and the Cross-Over and Mutation functions can be found in Appendix C.1.
3.3.2 ConfidenceTriggers-AutoDAN
As an alternative method for more naturalistic triggers, we propose a variant of ConfidenceTriggers based on AutoDAN [liu2024autodan] (ConfidenceTriggers-AutoDAN). AutoDAN is an algorithm for automatically generating stealthy jailbreak prompts through a hierarchical genetic algorithm. Featuring the same structure as ConfidenceTriggers, hierarchical perturbations are produced each generation at the sentence-level through synonym replacement based on scores from a momentum word score dictionary, and at the paragraph-level using GPT4 [OpenAI2024] to rephrase pairs of parent prompts for Cross-Over and Mutation. Full details and the algorithm are provided in Appendix D.
3.4 Design of Confidence Elicitation
Table 1: Description of CEMs.
CEMs Description Example Answer
Base Produce letter answer and confidence A, 80%
CoT Think step by step, produce letter answer and confidence at the end Since … therefore the answer is B, 80%
MS Break answers into multiple steps, produce confidence for each, produce answer and overall confidence Step 1: … 90%, Step 2, … 70%, …  Final Answer: C, 85%
SC Think step by step, produce letter answer and confidence at the end. Sample and average multiple answers. (1) … therefore the answer is B, 80% (2) I believe … so the answer is B, 90% (3) …, the answer must be B, 95%. Final Answer B, 88.33%
Our confidence elicitation methods are based on the framework proposed in [xiong2024can], where CEMs primarily consist of a prompting strategy (i.e., task instruction) for generating verbalized numeric confidence scores. CEMs can also include sampling strategies for generating multiple LLM responses, which are then aggregated to ideally produce a more representative confidence. We use the self-random sampling strategy, which samples $K$ different output sequences $\{\bm{\mathcal{Y}}_{1},\bm{\mathcal{Y}}_{2},\dots,\bm{\mathcal{Y}}_{K}\}$ using a specific temperature parameter, from which confidence estimates $\{\mathcal{C}_{1},\dots,\mathcal{C}_{K}\}$ are derived. To aggregate these disparate estimates, we take the average of the confidence estimates over the $K$ estimates as the final confidence estimate $C_{AVG}$, and choose the most common predicted label as the answer $\mathcal{A}_{FINAL}=\mbox{Mode}(\mathcal{A}_{1},\dots,\mathcal{A}_{K})$.
As described in Table 1, we propose to incorporate four main CEMs: Base, Chain-of-Thought (CoT), Multi-Step (MS), and Self Consistency Prompting (SC) [wang2023selfconsistency]. The first three methods are based on deterministic sampling. SC allows us to assess attack effectiveness under more random variation by averaging confidence across multiple sampled responses. Note that we utilize an additional form of SC called SC-CA (Self Consistency - Complete Average) in certain experiments, most notably for developing ConfidenceTriggers, since utilizing absolute model confidence is beneficial for optimizing the algorithm. While SC averages the confidence scores of only the majority voted answer out of K predicted samples, SC-CA is concerned with absolute confidence and hence averages the confidence scores of all K predicted samples irrespective of the majority voted class. This is useful in cases where priority is on confidence irrespective of label (please refer to Appendix B for the formulation of SC-CA). We also note that MS produces far less confident responses allowing us to ascertain attack performance in lower confidence scenarios. We focus on these CEMs since they are the most straightforward to attain and directly capture the model’s perception of confidence with minimal processing. Examples of prompts can be found in Appendix W.
4 Experiment Setup
Data.
We use popular Question Answering (QA) benchmark datasets that cover the generic, medical, and legal domains: (1) MedMCQA (MQA) [pmlr-v174-pal22a] (2) TruthfulQA (TQA) [lin2022truthfulqa] (3) StrategyQA (SQA) [geva2021strategyqa] We also use the AdvGLUE [wang2021adversarial] adversarial benchmark for confidence attack agnostic results, specifically the SST-2 [socher-etal-2013-recursive] set. Details can be found in Appendix C.
Models.
We focus on two LLMs of smaller and larger scale parameter sizes: Llama-3-8B [llama3modelcard] and GPT-3.5-turbo [OpenAI2024]. We also validate the results on GPT-4o [OpenAI2024] and Llama-3-70B [llama3modelcard]. Refer to Appendix C for details.
Evaluation Metrics.
The primary metrics are the average level of predicted confidence on the original (Cf.) versus adversarial examples (Adv. Cf.). We track what percentage of examples had attacks that successfully reduced the confidence (%Aff.), the average number of attack iterations for successful attacks (#Iters) to compare the attack efficiency, and the average level of confidence reduction on successfully attacked examples ($\Delta$ Aff. Cf.) i.e., ones that experience a confidence drop. We show which percentage of adversarial examples led to answer changes, both among originally correctly classified examples (%Flp(OC)) and incorrectly classified examples (%Flp(OW)).
Table 2: Effectiveness of different attack methods when targeting questions in QA prompts against a range of CEMs. Second column of Avg. shows the average distance between Cf. and Adv. Cf. Underlined are the most effective CEM results for each respective model/dataset/VCA method combination. In general, we can observe significant differences in Δ Aff. Cf. post attack (up to 70%) and high percentages of affected samples (up to 85%) from VCAs.
VCA-TF SubSwapRemove
Cf. Adv. Cf. %Aff. #Iters Δ Aff. Cf. %Flp(OC) %Flp(OW) Cf. Adv. Cf. %Aff. #Iters Δ Aff. Cf. %Flp(OC) %Flp(OW)
Llama-3-8B TQA Base 87.1 85.7 7.0 1.4 19.3 15.9 18.9 87.1 83.1 19.0 6.7 21.1 57.1 68.9
CoT 86.7 86.2 3.0 33.3 17.5 14.3 11.9 86.7 82.6 22.0 7.2 18.8 85.7 87.0
MS 61.0 51.8 43.0 11.8 21.4 33.0 30.3 61.0 49.8 52.0 7.5 21.6 42.9 53.2
SC 88.3 84.8 38.5 9.5 9.0 21.9 25.0 85.4 81.5 34.0 6.9 11.4 28.9 53.4
SC-CA 86.9 82.2 23.0 6.2 20.2 68.8 29.8 87.8 77.0 67.0 7.1 16.1 100.0 90.1
MQA Base 87.7 86.1 5.5 9.5 28.2 10.2 7.1 87.7 83.2 44.3 48.2 19.0 7.8 23.6
CoT 89.1 85.8 9.5 5.6 34.3 17.6 22.1 89.1 81.4 29.0 9.8 26.6 86.3 89.3
MS 58.1 44.8 53.0 16.4 25.2 50.5 52.7 58.1 44.3 55.5 9.8 24.9 46.8 67.0
SC 90.7 87.9 27.5 10.0 10.5 22.9 24.4 89.4 85.1 23.5 7.3 15.2 30.8 71.4
SC-CA 87.7 81.5 32.0 9.9 19.4 57.6 32.1 87.9 76.5 67.0 9.4 17.0 83.0 74.1
SQA Base 43.2 26.6 25.0 8.2 66.5 29.5 37.5 43.2 23.2 33.0 6.4 60.8 67.0 64.8
CoT 92.4 86.4 28.5 5.4 20.8 22.2 43.0 92.4 86.1 34.5 4.1 18.1 34.7 68.8
MS 83.6 66.2 65.0 11.4 26.7 33.1 35.9 83.6 67.4 68.0 6.2 23.8 50.0 42.2
SC 88.9 81.5 26.5 9.2 24.6 25.8 40.0 88.9 84.7 9.0 24.6 21.7 48.5 69.7
SC-CA 89.9 74.3 74.0 11.1 21.0 46.2 96.8 89.1 76.7 66.0 5.4 18.9 39.5 100.0
GPT-3.5 TQA Base 94.7 94.4 3.5 8.3 7.9 5.6 8.0 94.1 93.4 9.0 8.1 7.5 21.6 50.7
CoT 90.7 90.1 6.5 9.2 8.5 4.9 10.4 92.2 90.1 21.5 9.1 9.9 19.0 60.8
MS 69.5 62.4 44.5 12.9 15.9 17.4 30.6 68.8 58.7 58.5 7.2 17.2 28.1 72.3
SC 93.0 91.0 40.5 9.5 5.0 21.1 11.9 92.5 88.6 58.0 6.9 6.6 32.3 61.6
SC-CA 91.9 89.5 41.5 12.8 5.8 19.4 23.7 93.2 89.7 54.0 7.3 6.5 22.0 55.8
MQA Base 94.0 93.7 3.5 9.9 10.7 4.9 6.5 93.7 92.3 14.5 9.4 9.8 31.7 66.2
CoT 96.4 95.7 13.5 10.0 5.6 12.1 20.6 96.6 94.8 32.0 8.6 5.6 24.0 71.8
MS 62.5 53.3 51.0 15.2 18.1 23.1 50.0 63.6 52.9 60.0 8.7 17.8 43.2 77.0
SC 97.2 96.2 33.0 10.0 3.0 20.6 26.6 97.1 95.2 51.5 7.3 3.6 43.0 73.8
SC-CA 96.6 95.2 32.0 15.0 4.4 20.3 25.8 97.0 94.9 47.0 9.2 4.4 34.1 69.4
SQA Base 96.3 95.6 14.0 10.1 5.0 5.4 8.6 96.1 95.1 19.0 6.7 5.4 30.3 38.2
CoT 94.7 93.7 13.5 9.6 7.6 14.5 27.3 95.1 91.6 24.5 6.5 14.1 38.5 61.5
MS 85.2 71.1 59.0 11.8 24.0 25.2 45.6 85.5 66.9 70.5 5.9 26.4 49.7 60.8
SC 96.1 93.5 43.0 9.2 6.0 23.8 30.6 95.8 91.5 53.5 6.0 8.1 40.4 46.9
SC-CA 95.2 90.4 52.5 9.3 9.1 18.7 58.0 95.5 90.1 53.0 6.3 10.3 27.6 84.1
Avg. - 5.3 30.4 10.7 16.7 23.5 29.7 6.9 41.7 9.3 16.3 43.1 65.8
VCA-TB Typos
Cf. Adv. Cf. %Aff. #Iters Δ Aff. Cf. %Flp(OC) %Flp(OW) Cf. Adv. Cf. %Aff. #Iters Δ Aff. Cf. %Flp(OC) %Flp(OW)
Llama-3-8B TQA Base 87.1 84.2 10.5 5.1 27.1 15.9 18.9 87.1 80.8 25.5 22.8 24.7 73.0 89.2
CoT 86.7 86.0 4.0 5.8 16.9 28.6 10.4 86.7 78.9 27.0 26.5 29.1 100.0 97.9
MS 61.0 50.1 49.0 18.7 22.3 30.8 37.6 61.0 56.3 30.0 24.8 15.6 8.8 14.7
SC 87.9 82.6 39.5 14.2 13.2 35.4 25.7 85.5 83.7 12.0 25.0 15.2 11.9 24.2
SC-CA 88.2 81.5 36.0 10.3 18.5 60.7 44.2 87.9 81.6 54.0 25.0 11.6 96.2 74.7
MQA Base 87.7 85.6 7.0 17.4 28.9 13.6 6.3 87.7 83.2 22.5 25.0 19.7 75.0 80.4
CoT 89.1 84.8 12.0 8.9 36.1 19.6 25.5 89.1 84.3 23.0 25.0 20.9 98.0 96.0
MS 58.1 43.8 56.0 28.4 25.7 53.2 56.0 58.1 54.0 21.5 25.0 19.0 17.4 28.6
SC 90.6 85.5 30.5 14.9 16.5 32.5 34.9 90.0 89.0 6.0 25.0 17.9 15.7 22.4
SC-CA 86.3 79.4 42.5 17.4 16.1 68.9 46.8 86.8 79.5 62.0 24.8 11.8 60.0 56.6
SQA Base 43.2 24.1 27.5 15.0 69.5 42.0 50.0 43.2 20.3 37.0 24.7 62.0 75.9 71.6
CoT 92.4 84.2 31.5 8.9 26.0 34.7 49.2 92.4 85.1 38.0 24.0 19.0 59.7 78.1
MS 83.6 61.4 64.5 20.4 34.3 43.4 53.1 83.6 77.2 35.5 25.0 17.8 24.3 15.6
SC 91.1 77.2 71.5 12.8 19.6 29.3 30.2 91.7 88.8 27.0 24.6 10.8 15.1 11.1
SC-CA 88.0 69.1 78.5 19.2 24.2 61.3 100.0 87.4 74.3 75.0 24.5 17.5 20.7 100.0
GPT-3.5 TQA Base 94.6 94.4 3.5 13.7 6.4 5.5 8.3 94.4 93.4 13.0 25.0 7.5 33.3 59.5
CoT 91.5 90.7 9.5 17.2 8.4 4.3 20.2 90.3 87.2 30.5 24.2 10.2 34.5 71.6
MS 68.2 57.9 50.5 24.5 20.5 22.1 40.6 69.1 55.9 72.5 22.8 18.2 33.3 83.1
SC 93.3 90.9 46.0 14.2 5.3 31.5 30.1 93.6 92.5 26.0 25.0 4.2 8.3 19.1
SC-CA 92.9 89.5 46.5 19.4 7.4 23.8 32.1 93.5 90.0 71.5 24.7 4.8 22.7 42.6
MQA Base 93.9 93.4 5.0 8.4 10.0 4.9 6.5 93.9 92.1 18.5 24.5 9.7 62.2 77.8
CoT 97.2 96.2 18.0 15.8 5.6 10.2 20.6 96.5 94.7 33.0 25.0 5.6 42.4 80.9
MS 62.6 52.1 54.0 28.4 19.6 38.3 53.7 62.9 50.0 69.0 21.6 18.8 58.1 80.3
SC 96.6 95.3 33.5 14.9 4.0 26.8 22.6 96.8 95.2 47.0 25.0 3.5 57.0 83.1
SC-CA 96.8 94.9 37.5 23.3 5.0 28.4 25.8 97.3 95.1 50.5 24.8 4.4 40.4 65.6
SQA Base 96.1 95.3 15.5 17.2 5.0 6.8 16.4 96.2 95.1 23.5 25.0 5.0 31.8 27.9
CoT 94.9 93.6 16.5 13.1 8.2 15.1 29.6 95.0 90.4 32.0 24.8 14.3 51.0 63.3
MS 85.2 65.3 68.0 21.3 29.3 37.4 50.9 85.6 55.4 84.5 20.6 35.8 60.3 64.8
SC 95.9 92.2 51.5 12.8 7.2 25.8 42.9 96.3 89.1 69.5 24.6 10.3 47.0 46.9
SC-CA 95.6 89.3 57.0 19.6 11.1 19.0 66.0 95.8 87.2 79.0 24.9 10.9 17.5 71.7
Avg. - 6.9 35.8 16.0 18.3 29.0 35.2 6.5 40.5 24.4 15.9 45.1 60.0
5 Experiment Results
5.1 Confidence Agnostic Attack
We first determine whether CEMs are robust against generic confidence-agnostic adversarial attacks by comparing the confidences generated for the original instances in SST-2 [socher-etal-2013-recursive] and their AdvGLUE [wang2021adversarial] versions. The results are in Table 8 in Appendix E, where the difference in confidences are minor, with at most a 5% drop. CEMs are shown to be generally robust against confidence-agnostic perturbations, even though they are effective in reducing accuracy.
Figure 2: Averaged system prompt (Sys.) and demonstration attack (Demo.) performance across all models and the Base and CoT CEMs combined. Blue area over bars shows relative gains over equivalent query attacks while red show losses. We observe a high level of affected samples and significant average differences in confidence.
5.2 Perturbation-based Attack
To assess our proposed attack framework, we examine the four main VCAs against all of the CEMs across the three main datasets. For each dataset we randomly select 200 examples and test attacks against both correctly and incorrectly answered ones with the average performance across the examples. The results with user queries as the target are outlined in Table 2. We can see that the attacks are effective at reducing confidence on adversarial examples, yielding greater decreases in confidence scores compared to confidence-agnostic attacks. Nevertheless, the CEMs are largely resistant to drops in confidence of more than 50%, and the range of examples that experience a drop ranges from less than 5% up to 85%. When the attacks are effective, they lead to average confidence drops of 20%, which can have a great effect on human perception of veracity (consider a confidence of 90% versus 70%). The MS CEM, with far lower average confidence levels, is significantly more susceptible to the attacks in terms of confidence reduction, and in general the scenarios with lower average confidence are more susceptible, revealing that the greatest risk is for less confident approaches. A high percentage of the attack methods lead to answer changes, often flipping 50% up to 100% of answers, showing our framework leads to success in terms of the traditional adversarial objective.
In Figure 2 we present the averaged results across all datasets and models when using system prompts and demonstrations as attack vectors, focusing on the most robust CEMs in the Base and CoT methods (for full results, refer to Table 9 in Appendix F). System prompts and demonstrations are often more vulnerable to VCAs compared to queries, with the former being more effective at reducing confidence of the two, while demonstration-based attacks lead to higher rates of answer changes.
Table 3: Sub-sample of attacks using Llama-3-70B and GPT-4o (averaged across all datasets and CEMs). Best performance is underlined.
SubSwapRemove
Δ Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
Llama-3-70B Avg. 4.3 32.4 12.4 29.5 56.3
GPT-4o Avg. 5.7 48.0 11.4 27.4 53.6
VCA-TB
Llama-3-70B Avg. 3.8 26.6 12.8 14.9 28.4
GPT-4o Avg. 4.8 39.4 10.9 16.3 31.2
Results on GPT-4o and Llama-3-70B.
Due to the high cost of adversarial attacks, we design an experiment using the SOTA models, GPT-4o and Llama-3-70B, and run them over SSR and TB, covering the most effective of the different attack classes. The results (targeting user queries) are in Table 3 (full results in Table 10), confirming the findings in Table 2. In Table 11 we further test an additional three recently released models in Llama-3.1-8B [llama3modelcard], Llama-3.3-70B [llama3modelcard], and o3-mini [OpenAI2024] to show consistent VCA performance across recent architectures.
Table 4: The effectiveness of ConfidenceTriggers, ConfidenceTriggers-AutoDAN, and a random tokens baseline are in reducing downstream confidence scores. The highest averaged drops ConfidenceTriggers across all methods of confidence for a given model are in bold. The highest performances for each metric are underlined. Negative Δ’s show that average confidence increased post attack. We observe large Δ Cf. induced by optimizing a single set of tokens.
Random Tokens ConfidenceTriggers ConfidenceTriggers-AutoDAN
Δ  Cf. % Aff. ΔAff. Cf. %Flp(OC) %Flp(OW) Δ  Cf. % Aff. ΔAff. Cf. %Flp(OC) %Flp(OW) Δ Cf. % Aff. ΔAff. Cf. %Flp(OC) %Flp(OW)
Llama-3-8B TQA Base 0.9 17.0 5.0 13.8 40.7 3.6 18.8 19.1 16.7 33.6 0.4 16.5 1.4 0.6 15.6
CoT 1.3 33.0 4.0 81.8 71.4 29.5 66.0 44.7 100.0 67.7 1.9 15.0 6.7 1.3 30.6
MS 2.1 97.0 2.2 85.0 43.9 24.9 97.5 25.6 100.0 46.1 7.9 41.0 10.1 4.6 6.4
SC-CA 1.1 88.5 1.2 81.1 79.8 4.5 92.0 4.9 97.5 70.6 2.6 42.5 3.2 3.9 28.3
MQA Base 0.6 17.5 3.7 18.2 46.5 8.7 33.5 26.0 28.4 38.8 0.9 36.5 2.4 4.7 14.9
CoT -1.4 28.5 -4.8 56.3 52.0 9.9 39.0 25.3 67.4 72.0 0.4 25.5 1.7 16.2 26.5
MS 3.0 90.5 3.4 51.0 40.3 4.5 90.0 4.9 63.4 50.3 5.3 79.5 6.6 13.0 31.8
SC-CA -1.4 73.0 -1.9 62.2 76.1 3.6 84.5 4.3 82.5 70.0 1.5 64.5 2.3 18.3 44.0
SQA Base -6.3 27.0 -23.5 33.3 33.7 3.1 30.5 10.1 27.7 41.5 13.6 42.0 32.5 4.3 3.6
CoT 3.1 37.5 8.3 18.2 37.5 24.7 72.5 34.0 22.7 48.5 1.3 28.0 4.7 5.7 7.7
MS 1.9 91.5 2.1 24.7 25.2 6.2 81.0 7.6 23.4 34.8 3.0 61.5 4.9 5.6 9.2
SC-CA -0.9 91.0 -1.0 52.6 54.8 0.2 89.0 0.3 65.7 56.9 1.6 67.0 2.3 14.6 14.3
Avg. 0.3 57.7 -0.1 48.2 50.2 10.3 66.2 17.2 58.0 52.6 3.4 43.3 6.6 7.8 19.4
GPT-3.5 TQA Base -0.2 11.5 -1.7 4.7 8.3 0.2 15.5 1.5 6.3 23.6 9.0 68.5 13.2 1.6 33.3
CoT 0.7 25.5 2.7 7.9 20.5 0.2 15.5 1.5 5.0 27.8 4.4 67.0 6.6 6.5 32.5
MS 1.1 70.5 1.6 6.7 24.2 5.6 79.0 7.1 10.8 37.7 12.5 80.0 15.6 4.5 31.8
SC-CA 1.7 61.5 2.7 7.9 25.7 1.5 76.5 2.0 4.2 34.6 6.8 90.0 7.6 10.1 28.2
MQA Base 0.3 12.5 2.2 4.5 18.0 0.2 18.5 1.2 17.9 19.3 8.6 84.5 10.1 10.2 17.4
CoT 0.0 27.5 0.1 7.8 18.6 0.8 25.0 3.0 8.2 15.4 2.4 45.5 5.2 6.6 20.3
MS 2.6 65.0 4.0 10.8 31.4 3.4 63.0 5.4 8.0 13.8 7.8 75.0 10.4 3.8 19.1
SC-CA 0.9 53.0 1.7 4.4 29.2 1.0 46.5 2.1 6.1 20.9 2.2 67.0 3.3 11.3 22.4
SQA Base 0.7 15.5 4.4 3.9 6.8 1.3 23.5 5.5 9.9 7.6 6.1 70.0 8.8 8.2 14.1
CoT 0.4 20.0 1.8 9.0 18.2 0.1 21.5 0.5 11.3 8.5 2.1 34.0 6.1 10.3 14.1
MS 1.6 58.5 2.8 4.8 9.1 4.4 70.5 6.2 8.8 18.8 16.7 88.5 18.8 7.9 16.4
SC-CA 0.3 48.0 0.7 5.4 17.3 0.4 48.0 0.7 9.3 15.0 2.2 58.0 3.9 7.4 16.9
Avg. 0.8 39.1 1.9 6.5 19.0 1.6 41.9 3.1 8.8 20.2 6.7 69.0 9.1 7.4 22.2
5.3 Jailbreak-based Attacks
The performance of the optimized triggers (on examples not used to optimize triggers) in both the original and AutoDAN variation are shown in Table 4. Additionally, a random tokens baseline is included as a control. For this, we generate a prompt consisting of $M=20$ randomly selected confidence related words that are appended to the system prompt. Very high confidence drops (up to 29.5%) can be achieved with ConfidenceTriggers, although with inconsistency. ConfidenceTriggers-AutoDAN is less effective but can still lead to large $\Delta$ confidence (up to 16.7%). Both approaches outperform the random tokens baseline. ConfidenceTriggers also leads to a high percentage of answer changes, although the AutoDAN variant is less effective. Nevertheless, we show that triggers from the AutoDAN variant transfer across datasets in Appendix U.
5.4 Stability of Confidence
To understand the stability of verbal confidence, we track confidence and answers as words are removed in succession based on their importance score from user queries. This allows us to understand how confidence and answers change as the most critical words are removed, and when the input is completely stripped of its original contents. Figure 3 shows that the average confidence scores fluctuate inconsistently as tokens are removed. Even with most of the sequence missing, the difference largely remains below 15% compared to the original. This also demonstrates that reducing verbal confidence effectively requires specific optimization rather than degrading the input as much as possible. In contrast, answer changes follow a consistent pattern as they occur more often when the sequence is corrupted more. As a result, answer generation is more responsive to relevant factors in a sequence of text than confidence.
In Appendix G, we continue to study input corruptions by masking or randomizing confidence scores for each sub-step when using the MS method and observing how the LLM shifts its final predicted confidence when generating based on the corrupted input. The effects are found to be minor, with even the most extreme cases only leading to a drop of 30-40%. Consequently, we determine that generated confidence scores are stable with regards to serious text corruptions.
Figure 3: Average changes in confidence and answers as words are removed from queries according to their importance score, as a function of the percentage of word sequence is removed (Percent into Sequence) for Llama-3-8B and GPT-3.5. Positive $\Delta$ is for decreases in confidence compared to baseline and negative represents increases.
5.5 Confidence Behaviour
To understand the invariance of the tested LLMs, we provide the distribution of generated confidence scores in Appendix M, showing that LLMs rarely predict low confidences. We also determine that the most significant confidence changes occur on examples where the predicted answer changes (Appendix P), revealing how LLMs’ confidence is most vulnerable when they appear to be unsure of an answer. Nonetheless, in Appendix Q we test attack variants prioritized on only manipulating confidence and show that overall confidence vulnerability is not strongly tied to whether the model’s answer is ultimately changed due to adversarial perturbations.
We further evaluate attacks effectiveness when: (1) using a lower threshold for semantic similarity (Appendix K), and find that increases in attack effectiveness are minor, (2) LLMs are prompted to provide phrasal confidence from a Likert Scale (Appendix I) noting that attacks affect these disparate manners of confidence elicitation similarly, (3) using the Self-Probing CEM (Appendix J) we find that generating answers and confidence score in separate steps brings few benefits. Moreover, in Appendix H, when different phrases that ask LLMs to change their behaviour in terms of confidence are added to prompts, there are very large shifts in confidence. On the whole, these results reveal that LLMs’ maintain a high level of certainty irrespective of input, and behaviour primarily changes when explicitly instructed, though not often in ways that are expected.
Finally, we study alignment between numeric confidence scores and token log probability confidence in Appendix L, where we conclude that attacks create heightened misalignment in adversarial examples. This hurts model honesty by demonstrating that consistency between internal states and verbal outputs is degraded through minor perturbations. Amid the growing importance of calibration analysis, we observe how a notable degree of underconfidence is induced after attacks which aligns with our expectations (Appendix N). Despite generally high confidence levels, we highlight that LLM calibration does not entail honesty or lack thereof, since a model can become calibrated solely though imitation of a calibrated individual [zhou-etal-2023-navigating], hence we stress the importance of examining robustness and honesty regardless of inherent model calibration patterns. Nonetheless, in Appendix O we validate the calibration of our CEMs by presenting Brier scores [VERIFICATIONOFFORECASTSEXPRESSEDINTERMSOFPROBABILITY] across our results. The scores produced are reasonable (below those seen in works such as [galil2021disrupting]), and in most cases increase post attack showing that miscalibration is generally induced through VCAs. Lastly, we also compare AUROC scores prior to and after VCAs in Appendix V, where we find that AUROC decreases after attacks, confirming that VCAs are effective at degrading both confidence and predictive performance.
5.6 Defences Against Jailbreak-based VCAs
Table 5: ConfidenceTriggers-AutoDAN Paraphrase and SmoothLLM defense results against ConfidenceTriggers-AutoDAN on Llama-3-8B. Negative Δ’s show that average confidence increased post permutation.
SQA TQA MQA
Paraphrase
Δ Cf. O. Δ Cf. A. %Aff. %Adv. Aff. Δ Adv. Δ Cf. O. Δ Cf. A. %Aff. %Adv. Aff. Δ Adv. Δ Cf. O. Δ Cf. A. %Aff. %Adv. Aff. Δ Adv.
Base -24.8 -15.3 52.5 24.2 28.7 -1.6 7.2 34.0 42.4 23.2 0.5 34.5 44.0 84.8 44.0
CoT 1.4 1.7 44.0 8.9 27.5 5.2 24.0 43.5 47.8 53.2 7.9 11.0 39.0 36.9 30.7
MS -1.8 0.7 78.0 50.0 10.9 5.4 5.4 91.5 52.9 12.1 6.5 8.1 92.0 68.8 6.4
SmoothLLM
Base -14.4 -7.1 79.0 38.5 15.8 -1.6 0.1 72.5 37.0 5.8 5.2 5.3 71.5 54.5 3.6
CoT 5.8 8.9 93.5 87.0 6.0 3.8 4.7 95.0 67.5 6.2 11.4 10.2 94.5 86.0 15.4
MS -1.4 0.7 98.0 56.0 11.9 -1.7 -0.7 100.0 46.5 3.6 1.2 3.7 99.5 55.5 0.0
Jailbreak-based VCAs present a threat as although many detection systems for detecting and filtering general unsafe behaviour exist, none exist for detecting whether jailbreaking prompts are designed to influence confidence. To highlight the limitations in existing methods, we examine three typical categories of defences [xu-etal-2024-comprehensive]: two input permutation defences (Paraphrase defence [jain2023baselinedefensesadversarialattacks] and SmoothLLM [robey2023smoothllm]) and one self-processing perplexity filter defence [jain2023baselinedefensesadversarialattacks]. Refer to Appendix S for details. We experiment with jailbreak-based VCAs due to their more generalizable nature and many recent works are devoted to mitigation strategies against jailbreaks. We center our analysis on ConfidenceTriggers- AutoDAN, since the more naturalistic triggers offers greater difficulty to filter-based techniques compared to the main variant. Llama-3-8B serves as the target model. We use the same setup as for Table 4. Results on GPT-3.5 are in Appendix R.
Table 5 presents the results for the two input permutation defences. We show the difference in average confidence provided by the LLM between the original set of data and those with the permuted prompts without any trigger phrase (Δ Cf. O.) and with it (Δ Cf. A.). Additionally, we note what percentage of permuted examples with originally benign (% Aff.) and trigger-included prompts (% Adv. Aff.) had an altered generated confidence level compared to the original un-permuted version. Finally, we record the average decrease in confidence after input permutation on the trigger-included prompts (Δ Adv.). Our findings are that these defences lead to significant changes in confidence with both the inclusion and exclusion of a trigger phrase (differences up to 43%). With the majority of examples having their confidence unnecessarily altered from these defences, along with minimal decreases in attack effectiveness after adversarial prompt permutation, we deduce these defences are ineffective.
Table 23 in Appendix R presents the results of the perplexity filter and LLM-Guard [goyal2024llmguardguardingunsafellm], focusing on the percentage of user queries that experienced a confidence drop post VCA that get filtered, and which percentage of prompts with no trigger that would get filtered. In most cases for both methods the majority of these examples do not get filtered.
5.7 Perturbation-based VCA Defence Challenges
We detail how utilizing common defences against perturbation-based VCAs presents a major challenge. To illustrate the difficulty in applying a perplexity filter defence, in Table 6 we compare the average and standard deviation of the perplexity across normal, adversarial (generated using VCA-TF and VCA-TB), and a typical set of human generated text (500 samples from Twitter Sentiment140 [go2009twitter]) using Llama-3-8B on TQA. Through this, we show that despite the perplexity of perturbed inputs being higher than the base tested dataset, it still has a notably lower perplexity compared to real-world text obtained from social media which represents a typical range of informal and unrefined language that models can expect as input in many domains. This illustrates the difficulty of setting an effective perplexity filter without making strict assumptions about the input data distribution, as perplexity is not elevated beyond an unreasonable level.
Table 6: Using Llama-3-8B on TQA data, the average and standard deviation of the perplexity (Perp.) across original (O.), adversarial (Adv.), and data from Sentiment140 is presented.
O. Adv. Sentiment140
Avg. Std. Avg. Std. Avg. Std.
Perp. 70.3 84.1 350.1 640.2 456.7 1043.4
Given the lack of explicit mentions of confidence in perturbed adversarial examples, we demonstrate in Table 27 that by using a simple filter from prompting GPT-4.1 (detailed in Appendix T) to flag any inputs trying to manipulate confidence scores, it is possible to detect all direct confidence statements but none without them. This shows that while directly asking an LLM for a desired confidence level is an effective attack, it is very easily detectable compared to targeting confidence using indirect means like minor word or character level modifications, hence why they are the focus of our study.
6 Conclusion and Limitations
We have established the threat to LLMs from VCAs through a comprehensive study across different datasets, CEMs, and threat vectors. By formulating four perturbation-based and two jailbreak-based methods, we reveal serious vulnerabilities that can cause frequent and high levels of changes in verbal confidence and predicted answers. Through extensive analysis on the behaviour and properties of CEMs under a variety of scenarios and mitigation strategies, we demonstrate inadequacies with existing defences at fully mitigating the effects of attacks. Ultimately, as black-box LLMs become more widely adopted among both general users and in safety critical tasks where getting uncertainty estimates is key, we believe a higher focus needs to be ascribed not only to creating new and better calibrated black-box probability measures utilizing verbal confidence, but ensuring that such measures are also robust. This can be done by developing strategies for mitigating the ease with which they can be utilized for adversarial attacks and how simply they can be manipulated through trivial changes in input. In this sense, future methods need to be robust against VCAs but not so robust that they prevent LLMs in responding adequately to situations where the confidence should change.
The nature of running adversarial attacks requires many costly LLM queries, so we are unable to run more extensive experiments on larger language models. Nevertheless, we highlight two powerful LLMs at different scales, and corroborate our results using recent SOTA models. In addition, we primarily focus on character and word-level VCAs and have demonstrated their effectiveness. Sentence-level attacks can be constructed using a similar framework which we leave for future work. Our main VCA objective is also confidence reduction to mirror conventional adversarial attacks. VCAs aimed at increasing confidence are possible, but leave less room to study attack effectiveness due to the high confidence of victim LLMs. Lastly, we note that the datasets we test possess low ecological validity [devic2025calibrationcollaborationllmuncertainty]. Nevertheless, we have focused on well-studied datasets which are widely used in previous works. This is important, as due to the inherent variability of real-world data it is important to explore the idea of VCAs with well-tested datasets in tasks that previous work has covered for ease of interpretability. We strongly encourage future work to explore how the unique properties of different domains and tasks affect VCAs.
Appendix A Motivation
Industry Use of Verbal confidence
Firstly, we stress that studying verbal confidence and its robustness is an important topic given that there are numerous instances in industry where systems have been developed that utilize some form of LLM verbal confidence in their methods to conduct confidence estimation. They utilize verbal confidence largely due to limitations with black-box API access and since the scores are not arbitrary [pmlr-v235-mohri24a, Omar2024.08.11.24311810] they have value as a confidence measure. These use-cases are shown to ultimately be concerned about robustness [chen-mueller-2024-quantifying, macdonald2024generative, zhang2023pacelmpromptingaugmentationcalibrated] which we will highlight. chen-mueller-2024-quantifying develop a Trustworthy Language Model (TLM) where Reliability and explainability is added to every LLM output and decision-making is conducted using trustworthiness scores. This score quantifies the certainty of the model to its answers by combining checking the consistency and the level of verbal numeric confidence of the model’s responses. zhang2023pacelmpromptingaugmentationcalibrated, in the context of cloud incident root cause analysis, develop a method for on-call engineers to make decisions on whether to adopt model predictions using a method that incorporates an LLM-based confidence estimator using verbal confidence due to the black-box nature of the APIs of SOTA models. In macdonald2024generative they build an LLM-based product to structure unstructured data and score customer conversations for developing sales and customer support teams. These scores are based on level of customer satisfaction and given relevant (proprietary) context, the LLM produces a score (verbally) and reasoning for that score. In all of these use-cases it is clear that compromising the robustness of these verbal scores creates a risk of system malfunction or a degradation in operation. Likewise, in pmlr-v235-mohri24a, although not directly related to industry, show the usefulness of utilizing verbal confidence in a specific application, which is that of generating conformal factuality guarantees for language models. They reveal that verbal confidence outperforms the arbitrary confidence baselines (i.e., randomly generated confidence scores and ordinal confidence), demonstrating its value.
Importance of Calibrated Generated Confidence Scores in Real World Applications
Many prior works such as dhuliawala-etal-2023-diachronic, 10.1145/3351095.3372852, 10.1145/3491102.3501967, have established how harmful miscalibrated (skewed) confidence scores are in human studies and downstream human-AI collaboration applications, showing that user behaviour changes greatly in response to AI confidence scores, leading to over reliance when an AI system is overconfident, as one example. Most notably, dhuliawala-etal-2023-diachronic shows in the context of human collaboration with LLMs how un-confidently correct stimuli reduce user trust with only a few miscalibrated instances creating long-term poor performance on tasks.
Moreover, works such as bernsohn-etal-2024-legallens, 50013, 10399826, 10.1136/amiajnl-2011-000291 highlight the impacts of altered confidence scores on downstream authority mechanisms in legal and medical domains, since many such applications that involve tasks such as diagnosis rely on model uncertainty to determine further review. In bernsohn-etal-2024-legallens as an example, in the task of detecting legal violations, entities with high confidence undergo additional processing, and therefore reducing the confidence estimates in the examples would compromise the operation of the system.
Risk of Adversarial Attacks Based on Confidence
In addition to harm done by reduced confidence scores (when answers remain unchanged), our key contribution is also to demonstrate verbal confidence can be used to effectively craft conventional adversarial examples (i.e., ones that lead to answer changes) as well, which presents issues in practical scenarios (e.g., misclassified cases). This is important because such scores are easy to obtain for adversaries for black-box models, and can be used to manipulate the confidence measures that can most easily be asked for and understood by human users.
Without our work, it is not clear how effective the proposed attacks would be, given the very different properties of using generated outputs versus internal model states or a proxy white-box model. If adversaries are able to easily generate attacks using solely the generated output, this presents a great security concern.
Our formulation of this issue also has contributions in terms of its implications towards model honesty and how models express their level of knowledge/competence through confidence. A key aspect of this study is to determine how easy their behaviour is to manipulate through various types of mostly superfluous modifications, and through our analysis determining among which aspects there are the most vulnerabilities. In short, one of our objectives is to see how models “perceive” confidence overall and hence we look at multiple different aspects of how models react to different changes of input. We primarily test generalist models to get a broad view of the problem of robustness and to avoid making any conclusions that would only apply to a singular domain.
Threat Model
Similar to established adversarial attack setups, users of the LLM do not necessarily have to be the ones conducting the VCAs. Long-tailed examples can naturally happen to jeopardize the victim model including variations (like typos) in their input that can affect confidence and one aspect of this analysis is to see how easy it is to change confidence scores like this.
Secondly, our overall threat model is no different compared to other NLP adversarial attack works; adversaries can craft adversarial examples independently without being the direct user of the system being attacked. If an adversary is able to corrupt a series of queries beforehand, then when somebody (i.e., a user) tries to evaluate these queries asynchronously they get miscalibrated confidence scores without being aware of any changes. For example, samples from a database can be manipulated by an adversary and swapped in without a user being aware.
When we refer to attacks against user queries, we do not necessarily mean the current user is intentionally crafting an example to see how much they can lower the model confidence since the model can simply be prompted to change the score (as we show can be effectively done in Appendix H). Nonetheless, various changes the user obliviously makes to their query can also affect the confidence hence we study the impact of perturbations and corruptions. Ultimately, we wish to see how effectively it can be done by a different source such as an external adversary.
We study three different threat vectors. An adversary can also target demonstrations or the system prompt too without a user being aware when they send their own unmodified query. This can be a form of data poisoning attack since, for example, adversarially modified demonstrations could be used by a developer of an LLM system unknowingly from a database of seemingly benign demonstrations.
Real-World Use Case
A hypothetical scenario for the threat of VCAs could be that of a legal professional using a proprietary LLM to determine whether a series of prior cases are relevant to their current case, where highly relevant/high confidence predictions above a threshold are used to tell them specifically to further examine the case. An adversary wishing to target this legal professional could attack the database of cases so that a series of low confidence predictions are produced on attacked examples, potentially causing the professional to miss relevant cases when presenting them to the LLM, leading the model to cease being useful, creating severe inefficiencies and ultimately eroding user trust.
Appendix B Details of Perturbation-based Attacks
The general algorithms for the four different character and word-level adversarial attack methods we cover can be seen in Algorithm 2. For VCA-TF and VCA-TB we make modifications on the black-box algorithms to make the algorithm work in the context of attacking confidence scores generated using CEMs, as these algorithms are originally based on using the predicted softmax class-wise probabilities derived from model logits. Refer to more details in DBLP:conf/aaai/JinJZS20 and DBLP:conf/ndss/LiJDLW19 respectively for more detailed settings as we are faithful to the original algorithm except with respect to the scoring functions used. For the Typos and SubSwapRemove attacks, we make use of our own design choices which we detail below. In the algorithm, the potential options for the target attack text $\mathbf{X}$ are user queries, system prompts, or demonstrations. Additional input text $\mathbf{Q}$ refers to the two remaining text inputs that are not undergoing the attack (e.g., if the user query is targeted, $\mathbf{Q}$ represents the system prompt and demonstration). The parts that are attacked for user queries refer to the question only, without the multiple choice options. For system prompts the entire prompt detailing the task instructions (i.e., how to generate answer and confidence score) can be attacked. For the demonstration attacks, the demonstration is a one-shot example for a given CEM. Within the example, the question, multiple choice options, and the expected answer can be targeted.
For VCA-TB, the possible random bugs that can be generated include inserting spaces into words, deleting characters, swapping adjacent letters, swapping characters with visually similar characters, and substituting words with one of its nearest neighbours. We use WordNet [wordnetcitation] to search for synonyms for each token and calculate the semantic similar using Euclidean distance with GloVe embeddings [pennington-etal-2014-glove]. Likewise, with VCA-TF, we use the same approach to generate a list of synonyms, although we also ensure that synonyms contain the same part of speech (POS) tags using NLTK [https://doi.org/10.48550/arxiv.cs/0205028].
To perform the Typos attack, we generate a series of common typos in a sentence (GenerateCommonTypos function). These typos include substitutions with nearby characters on an English keyboard, random character deletions, and adjacent character swaps. A probability of $0.1$ is used to select the percentage of characters in a full text sequence to modify each iteration. We cap the number of iterations at 25.
In the SubSwapRemove algorithm, MOD function applies a random perturbation to a given token. This modification can be either a synonym substitution, a swap with an adjacent token, or the token is deleted from the sequence. The probability of selecting each modification is equal. For the synonym substitution, we sample five top synonyms similar to VCA-TF and keep the substitution that results in the greatest drop in confidence if a drop occurs at all with any of the synonyms.
Across all attacks, we stop attacks once the original predicted answer is changed and record this along with the final confidence to keep track of how easily the answer can change.
The description for the SC-CA CEM and an example is provided in Table 7.
Algorithm 2 Overview of Tested Attack Methods
1:Input: Confidence Estimation Method $z$ for LLM, Tokenized Target Text $\mathbf{X}$, Additional Input Text $\mathbf{Q}$ Sentence Similarity Function SIM(), Perturbation Similarity Rejection Threshold $\epsilon$, Max Iterations $\tau$, # Synonym Candidates $N_{syn}$
2:Output: Adversarial Example $\hat{\mathbf{X}}$
[…]
5:
6:Importance Score Based Methods:
7:Filter Stop Words
8:for all $\hat{x}_{i}\in\mathbf{\hat{X}}$ do
[…]
10:end for
[…]
12:Sort $\mathbf{W}$ according to $\mathbf{J}$
13: -VCA-TB-
14:for all $w_{i}\in\mathbf{W}$ do
15:  BugBug $\leftarrow$ Randomly Selected TextBugger Bug on $w_{i}$
16:  $\mathbf{\hat{X}}_{temp}$ $\leftarrow$ $\mathbf{\hat{X}}_{w_{i}<>Bug}$ $\triangleright$ Bug replaces original token in $\mathbf{\hat{X}}$
17:  if SIM($\mathbf{X},\mathbf{\hat{X}}_{temp})>\epsilon$ then
18:   if $z(\mathbf{Q},\mathbf{\hat{X}}_{temp})<\mathcal{C}$ then
19:     $\mathbf{\hat{X}} \leftarrow$ $\mathbf{\hat{X}}_{temp}$
20:   end if
21:  end if
22:end for
23: -VCA-TF-
24:for all $w_{i}\in\mathbf{W}$ do
25:  Candidates $\leftarrow$ Generate top $N_{syn}$ POS tag filtered candidate synonyms for $w_{i}$
26:  topscore = null
27:  for all $\mbox{cnd}\in\mbox{Candidates}$ do
28:   $\mathbf{\hat{X}}_{temp}$ $\leftarrow$ $\mathbf{\hat{X}}_{w_{i}<>cnd}$ $\triangleright$ cnd replaces $w_{i}$ in $\mathbf{\hat{X}}$
29:   if SIM($\mathbf{X},\mathbf{\hat{X}}_{temp})>\epsilon$ then
30:     if (topscore = null AND $z(\mathbf{Q},\mathbf{\hat{X}}_{temp})<\mathcal{C}$) OR ($z(\mathbf{Q},\mathbf{\hat{X}}_{temp})<\mbox{topscore}$) then
31:      $\mathbf{\hat{X}} \leftarrow$ $\mathbf{\hat{X}}_{temp}$
32:      topscore == $z(\mathbf{Q},\mathbf{\hat{X}}_{temp})$
33:     end if
34:   end if
35:  end for
36:end for
37:
38:Typos:
39:for $\tau$ iterations do
[…]
41:  if SIM($\mathbf{X},\mathbf{\hat{X}}_{temp})>\epsilon$ then
42:   if $z(\mathbf{Q},\mathbf{\hat{X}}_{temp})<\mathcal{C}$ then
43:     $\mathbf{\hat{X}} \leftarrow$ $\mathbf{\hat{X}}_{temp}$
44:   end if
45:  end if
46:end for
47:
48:SubSwapRemove:
49:Filter Stop Words in $\mathbf{\hat{X}}$
50:for all $\hat{x}_{i}\in\mathbf{\hat{X}}$ do
51:  MOD $\leftarrow$ Randomly Choose Perturbation Type from {Sub, Swap, Remove}
[…]
53:  if SIM($\mathbf{X},\mathbf{\hat{X}}_{temp})>\epsilon$ then
54:   if $z(\mathbf{Q},\mathbf{\hat{X}}_{temp})<\mathcal{C}$ then
55:     $\mathbf{\hat{X}} \leftarrow$ $\mathbf{\hat{X}}_{temp}$
56:   end if
57:  end if
58:end for
Appendix C Detailed Experimental Settings
Dataset Details: (1) MedMCQA [pmlr-v174-pal22a] is a multiple-choice question answering (MCQA) dataset consisting of 2816 questions with four choices that are based on medical entrance exam questions. (2) TruthfulQA [lin2022truthfulqa] is a MCQA dataset consisting of questions that cover a broad range of safety-critical topics like health, law, finance and politics and is designed so that humans can easily answer incorrectly due to misconceptions. We use the MC1 (Single-true) subtask, and only use multiple choice questions with 4 total answers for consistency yielding 203 total examples. (3) StrategyQA [geva2021strategyqa] consists of 2,780 true or false open-domain, general topic questions that require implicit reasoning steps to answer. Lastly, we use SST-2 [socher-etal-2013-recursive] from the popular AdvGLUE [wang2021adversarial] adversarial benchmark for our baseline confidence attack agnostic results. In the benchmark, SST-2 consists of 137 movie reviews annotated for binary (positive/negative) sentiment. Instances in AdvGLUE underwent a comprehensive series of word-level, sentence-level, adversarial algorithmic modifications, in addition to human-generated perturbations to most effectively deceive a wide range of language models. For our experiments, we use the development sets for MedMCQA and TruthfulQA, and the training set for StrategyQA.
Model Details: Llama-3-8B [llama3modelcard] and GPT3.5-turbo [OpenAI2024] are primarily used for evaluation, with GPT-4o [OpenAI2024] and Llama-3-70B [llama3modelcard] being utilized for certain additional results. The Llama models are ran locally while the GPT models are ran through the OpenAI API. In terms of sampling parameters for each model, we utilize a temperature of 0.0 for Base, CoT, and MS method, along with a top_p of 0.95. For SC and SC-CA, we set the temperature to 0.7 and sample three different responses and average the confidence scores. We limit the number of generated tokens to 400. For the MS method, we calculate the geometric mean of the generated confidence scores for each step including the final.
Sentence Similarity: To ensure adversarial examples preserve the original meaning, we encode both the original and adversarial example using Universal Sentence Encoder (USE) [cer2018universal]. We then calculate cosine similarity between the vectors and forgo perturbations that cause a fall below a threshold of 80% similarity to keep adversarial examples retaining similar meaning of the original sentences similar to the approach in wang2023adversarial.
Confidence Extraction: Answers and confidence are extracted from the generated answer of the LLM using robust regular expressions looking for all possible mentions of letter answers and their corresponding numeric confidence. In the rare cases where models refuse to answer or provide a confidence score (or the regular expression fails), the final answer defaults to none-of-the-above and the final confidence score is set to be one with max entropy possible (i.e., $1/(\mbox{Number of Multiple Choice Options})$) respectively. Over the course of optimizing an attack, if no confidence is provided, the perturbation is rejected unless the answer choice flips as well (signifying a successful attack).
Table 7: Description of SC-CA CEM.
CEMs Description Example Answer
SC-CA Think step by step, produce letter answer and confidence at the end. Sample and average multiple answers. (1) … therefore the answer is B, 90% (2) I believe … so the answer is B, 95% (3) …, the answer must be A, 80%. Final Answer B, 88.33%
C.1 ConfidenceTriggers Settings
To optimize the performance of ConfidenceTriggers, we evaluated a variety of algorithm parameters. However, given the significant amount of iterations necessary to optimize the algorithm, we limited some parameters like trigger prompt length while still aiming for strong results. Regarding our training text corpus, for MedMCQA and StrategyQA, we use 70% of the data for optimizing the triggers, and evaluate it on the remaining 30% (out of which 200 random samples are selected). Given the relatively smaller size of TruthfulQA, we train the triggers on questions with five multiple choices options (yielding 197 different examples) and evaluate on the 203 questions with four options. For our trigger prompt length $\beta$ we choose 20 tokens, and for the total number of prompts we attempt to optimize $\alpha$ we also choose 20. We experimented with optimizing between 20 and 50 total generations. For the number of samples for the initial estimation $S$, we choose 200, and the value $\xi$ for each iteration we choose 12, leading to 48 different sampled training cases used for estimation for each prompt each generation. We choose a fifth of prompts (i.e., 4) with the lowest loss to be elites. An example of the generated triggers can be found in Appendix X.
The InitializePrompts($\alpha,\beta$) function in the main algorithm that is used to create the initial set of $\alpha$ triggers before optimization occurs, essentially for each trigger generates a string of a $\beta$ number of random tokens that are uniformly sampled from the set of confidence related words. The AverageConfidence function is defined in Algorithm 5.
Our procedure for Cross-Over and Mutation functions remains largely faithful to lapid2023open and are standard implementations for these functions given two parent sequences of data. Cross-Over is the standard Single point Cross-Over [10.7551/mitpress/1090.001.0001] operation and Mutation is Random (uniform) Mutation [10.5555/534133] where words at randomly chosen indices are replaced. Pseudo code for Cross-Over and Mutation can be found in Algorithm 3 and 4 respectively. The tournament select process is top-K like in 10.1162/evco.1996.4.4.361.
Algorithm 3 ConfidenceTriggers Cross-Over
1:Input: Initial Parent Prompts $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$
2:Output: Child Prompts $\mathbf{C}_{1}$ and $\mathbf{C}_{2}$
3:SplitPoint = Choose random integer between 0 and m-2
[…]
Algorithm 4 ConfidenceTriggers Mutation
1:Input: Initial Parent Prompts $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ of length $\beta$
2:Output: Child Prompt $\mathbf{C}$
3:RandomOption $\leftarrow$ Choose either $\mathbf{P}_{1}$ or $\mathbf{P}_{2}$ to mutate randomly
4:IndicesToFlip $\leftarrow$ Sample $20\%$ of indices in $\mathbf{P}_{1}$ to mutate
5:$C=[]$
6:for j = 0,…, $\beta$ do
7:  if j in IndicesToFlip then
8:    Add random word to $\mathbf{C}$
9:  else
[…]
12:    else
13:     Append $\mathbf{P}_{2}[j]$ to $\mathbf{C}$
14:    end if
15:  end if
16:end for
Algorithm 5 ConfidenceTriggers AverageConfidence Function
1:Input: Confidence Estimation Method $z$, # Samples for Estimation $n$, Prompt Set $\mathbf{P}$
2:Optional: Trigger string $t$
3:Output: Average confidence $C$
[…]
5:if $t$ then
6:  $\mathbf{Confidences_{n}}\leftarrow$ Generate confidence using $z$ for $p\in\mathbf{P_{n}}$ with $t$ appended to $p$
7:else
8:  $\mathbf{Confidences_{n}}\leftarrow$ Generate confidence using $z$ for $p\in\mathbf{P_{n}}$
9:end if
[…]
Appendix D ConfidenceTriggers-AutoDAN Details
In this section, we present the algorithm along with the details for the AutoDAN variation of the ConfidenceTriggers algorithm. The general algorithm can be found in Algorithm 6 and generally follows the original version of ConfidenceTriggers but is modified to incorporate the steps and methods of the AutoDAN-HGA algorithm in liu2024autodan, in that we make use of the hierarchal structure of prompts and incorporate both sentence and paragraph level optimizations.
A few key differences exist between ConfidenceTriggers-AutoDAN and AutoDAN-HGA due to the nature of attacking confidence compared to jailbreaking. Firstly, rather than the initial adversarial prompts being based on previously successful manually curated jailbreak prompts, we generate a series of sentences using GPT4 that indirectly encourage a model to output low confidence scores. We limit the initial prompts to only a single sentence and not paragraph length like in the original variant for simplicity given that pre-established multi-sentence length triggers for attacking confidence are not available. Other aspects of the algorithm, like the functions for Cross-Over and Mutation are largely faithfully adapted, but with some changes. Each of the main functions in the algorithm are detailed below.
InitializePrompts($\alpha$) function generates $\alpha$ different prompts using a proxy model (GPT4) that encourage the production of low confidence scores (Diversification). The model fed a prompt stating to “generate $\alpha$ sentences that would encourage a model to produce low confidence scores in a machine learning context but in an indirect manner.”
Momentum Word Score is calculated the same way as in AutoDAN-HGA. A dictionary is established in which words from each of the prompts are added if they are not part of a collection of common stop words, punctuation, proper nouns etc. When a word is added as a key its score is added as a value. The score is derived from the loss of the prompt it originally came from. If the same word is first being added and it is present in multiple prompts, the score is averaged. If a word has a key it is updated using a momentum mechanism, whereby the updated value is averaged between its current (previous) value and the newly computed $(\mbox{current value}+\mbox{new value})/2$. The words in the dictionary are finally ranked based on their scores in descending order.
Replace Words with Synonyms functions by looping through each word in a prompt, and a series of synonyms are found for each. The word dictionary is searched using these synonyms and if one is present then it probabilistically replaces the original word based on the relative scores of all the other synonyms that are present in the dictionary.
The process of selecting parent prompts for Cross-Over and Mutation involves weighted random sampling. $\alpha-E$ prompts are randomly sampled without replacement with the weights based on the loss i.e., more effective prompts are more likely to be sampled. Each prompt is added to a list and then is paired up with its neighbour in the list (the first and second, third and fourth …). If there is an odd number of sampled prompts, then the last and the first in the list are paired again. The sampled prompts then undergo Cross-Over and Mutation.
For Cross-Over, if a randomly sampled float between 0 and 1 is above the threshold, then the process occurs for a pair of prompts. In our experiments, we use a threshold of 0.5. As prompts do not span multiple sentences, performing cross-over by splicing together different sentences from both prompts as in the original algorithm would not work. Instead, Cross-Over works by using a powerful LLM (GPT4) to combine the two parent prompts as succinctly as possible. The following system prompt is fed to the model: “You are a helpful and creative assistant who writes well.” Afterwards, the user prompt is as follows “Please combine the contents of the following two sentences (prioritizing preserving the first sentence) while trying to keep the length as minimized as possible (below 100 words). Please only output the revised version.” By feeding the two prompts but in different orders, this yields two child prompts with combined content but with different orders/priority.
The Mutation function occurs in all cases. The two parent prompts are fed individually into a powerful LLM (GPT4) with the same system prompt as Cross-Over but with the following user prompt: “Please revise the following sentence with no change to its length and only output the revised version.” This induces a degree of variance into each child prompt.
A total of 100 generations is used to optimize the prompts. The remaining settings are consistent with those detailed in Appendix C.1. An example of the final generated triggers can be found in Appendix X.
Algorithm 6 ConfidenceTriggers-AutoDAN
1:Input: Set of confidence estimation methods $\mathbf{Z}$, training text corpus $\mathbf{X}$, # generations $G$, # prompts $\alpha$, # initial samples for estimation $S$, # training samples per iteration $\xi$, # elites $E$
2:Output: Trigger String $\mathbf{T}$
3:for all $z\in\mathbf{Z}$ do
4:  Random subset sample $S$ cases from $\mathbf{X}$
5:  $\mathcal{L}_{z}$ $\leftarrow$ AverageConfidence($z$, $S$) $\triangleright$ Base confidence w/o any trigger
6:end for
7:$P$ $\leftarrow$ InitializePrompts($\alpha$)
8:for $g=1$ to $G$ do
9:  for all $p\in$ $\mathbf{P}$ do $\triangleright$ Evaluate fitness of each prompt
10:    for all $z\in\mathbf{Z}$ do
11:     Random subset sample $\xi$ examples from $\mathbf{X}$
12:     $\mathcal{L}_{p_{g}}$ $\leftarrow\mathcal{L}_{p_{g}}\mathbin{\|}$ AverageConfidence($p$, $z$, $\xi$) - $\mathcal{L}_{z}$
13:    end for
14:    ${L}_{p_{g}}=\sum\mathcal{L}_{p_{g}}/|\mathbf{Z}|$
15:  end for
16:  Sort $\mathbf{P}$ by $\mathcal{L}_{p_{g}}$
17:  $\mathbf{P}_{g}$ $\leftarrow$ [][\;]
18:  Add $E$ lowest loss prompts $p$ to $\mathbf{P}_{g}$ as elites
19:  Sentence-Level:
20:  Calculate momentum word score
21:  Replace words with synonyms for each $p\in$ $\mathbf{P}$
22:  Paragraph-Level:
23:  Use weighted random sampling to select $\alpha-E$
24:    parent prompts into a list $\mathbf{P}_{parents}$
25:  Pair each $p_{ADV}$ in $\mathbf{P}_{parents}$ together
26:  for all $(p_{a},p_{b})$ unique pairs in $P_{parents}$ do
27:    rand $\leftarrow$ Random float between 0 and 1
28:    if $\mbox{rand}<0.5$ then
29:     $p_{a}$, $p_{a}$ $\leftarrow$ CrossOver($p_{a}$$p_{b}$)
30:    end if
31:    $p_{a}$ $\leftarrow$ Mutation($p_{a}$)
32:    $p_{b}$ $\leftarrow$ Mutation($p_{b}$)
33:    Add $p_{a}$, $p_{a}$ to $\mathbf{P}_{g}$
34:  end for
35:  if $|\mathbf{P}_{g}|$ >> $|\mathbf{P}|$ then
36:    Keep first $|\mathbf{P}|$ prompts in $\mathbf{P}_{g}$
37:  end if
38:  $\mathbf{P}$ $\leftarrow$ $\mathbf{P}_{g}$
39:end for
40:for all $p\in$ $\bf{P}$ do
41:  for all $z\in\mathbf{Z}$ do
42:    Random subset sample $S$ cases from $\mathbf{X}$
43:    $\mathcal{L}_{p_{G}}$ $\leftarrow\mathcal{L}_{p_{G}}\mathbin{\|}$ AverageConfidence($p$,$z$,$S$)
44:  end for
45:  $\mathcal{L}_{p_{G}}\leftarrow$ $\sum\mathcal{L}_{p_{G}}$ / $|Z|$
46:end for
[…]
Appendix E Confidence Agnostic Attacks
Table 8: Results of confidence agnostic attacks on SST2. Difference in average confidence and accuracy between unmodified and adverserial instances are reported, along with the percentage of examples being affected.
Cf. Adv. Cf. Acc. Adv. Acc. % Aff.
Llama-3-8B Base 74.0 71.0 61.8 42.0 46.6
CoT 88.2 87.2 44.3 42.0 72.5
MS 80.2 81.3 49.0 48.1 83.2
SC-CA 74.6 69.6 51.9 51.9 97.7
GPT-3.5 Base 92.7 90.4 100.0 86.3 45.0
CoT 93.5 89.7 100.0 87.0 39.7
MS 90.3 85.6 100.0 86.3 75.6
SC-CA 94.6 90.7 100.0 85.5 64.1
We examine if CEMs are robust against generic confidence-agnostic adversarial attacks. To this end, we compare the confidences generated for the original instances in SST-2 [socher-etal-2013-recursive] and their AdvGLUE [wang2021adversarial] versions since these comprehensive adversarial modifications were solely optimized for misclassifications. The results are in Table 8, where the difference in confidences are minor, with at most a 5% drop despite a high percentage of samples being effected. CEMs are shown to be generally robust against confidence-agnostic perturbations, even though they are effective in reducing accuracy by 15-20%.
Appendix F Full Results Tables
The full results for the system prompt and demonstration attacks are in Table 9. Likwise, the full results for attacks against Llama-3-70B and GPT-4o are in Table 10.
Table 9: Effectiveness of confidence adversarial attacks against system prompts or one-shot demonstrations.
System Prompt Demonstration
VCA-TB
Cf. Adv. Cf. % Aff. #Iters ΔAff. Cf. %Flp(OC) %Flp(OW) Cf. Adv. Cf. % Aff. #Iters ΔAff. Cf. %Flp(OC) %Flp(OW)
Llama-3-8B TQA Base 87.1 84.7 11.0 42.9 21.6 8.7 16.2 87.1 80.1 27.0 11.8 26.0 32.5 32.4
CoT 86.7 81.2 18.0 22.3 30.7 71.4 46.6 86.7 81.1 20.5 19.9 27.1 42.9 45.6
MQA Base 87.7 82.9 20.5 37.2 23.5 9.1 21.4 87.7 83.5 41.5 10.8 10.1 28.4 42.0
CoT 89.1 83.1 13.5 31.7 44.4 29.4 32.2 89.1 81.7 35.0 14.5 21.3 45.1 58.4
SQA Base 43.2 36.5 17.0 39.1 39.8 25.0 25.0 43.2 33.5 28.5 24.7 34.3 66.1 64.8
CoT 92.4 60.0 55.0 32.9 58.9 56.9 82.0 92.4 80.8 50.0 26.3 23.1 36.1 75.0
GPT-3.5 TQA Base 94.7 92.7 21.5 80.3 9.7 15.6 11.1 94.7 93.2 19.5 12.6 7.7 9.6 18.7
CoT 92.1 89.2 28.0 127.3 10.5 21.1 24.7 92.2 90.0 21.0 23.9 10.5 26.9 32.1
MQA Base 93.8 91.6 22.0 60.0 9.9 11.7 12.5 93.9 92.3 21.0 17.5 7.5 15.7 23.5
CoT 96.3 94.2 35.0 116.9 6.0 15.6 30.8 97.0 95.5 27.0 35.7 5.6 15.1 40.5
SQA Base 96.4 95.3 22.0 44.0 5.0 0.8 2.9 96.3 95.2 21.5 33.5 5.1 4.5 14.9
CoT 95.8 92.9 28.5 39.0 10.2 9.9 22.9 95.0 94.3 11.5 35.3 6.5 12.0 38.0
Avg. - 5.9 24.3 56.1 22.5 22.9 27.4 - 4.5 27.0 22.2 15.4 27.9 40.5
VCA-TB
Llama-3-8B TQA Base 87.1 84.4 13.5 40.0 20.2 9.5 20.3 87.1 81.0 26.5 20.0 22.9 34.9 35.1
CoT 86.5 82.2 14.4 31.4 29.3 75.0 43.0 86.7 78.9 21.5 22.8 36.4 42.9 45.6
MQA Base 87.7 83.7 20.5 44.2 19.6 5.7 23.2 87.7 83.2 43.0 16.2 10.4 31.8 43.8
CoT 89.1 82.5 14.5 30.6 45.6 33.3 32.2 89.1 80.7 38.0 17.6 22.1 56.9 59.7
SQA Base 43.2 35.4 20.0 39.0 38.9 29.5 29.5 43.2 32.4 28.5 27.2 38.2 63.4 72.7
CoT 92.4 56.0 56.0 36.8 65.0 52.8 82.0 92.4 79.9 50.0 29.6 24.9 38.9 75.0
GPT-3.5 TQA Base 94.7 92.3 25.0 166.7 9.8 15.7 11.0 94.8 93.6 19.0 17.7 6.6 9.4 8.2
CoT 91.3 89.1 26.5 207.3 8.4 18.6 24.4 92.3 89.6 28.5 31.7 9.3 32.3 35.5
MQA Base 93.9 91.3 26.0 106.0 9.7 12.5 16.3 93.8 91.9 19.5 26.6 9.5 21.2 24.4
CoT 96.8 94.8 35.5 45.5 5.7 17.0 30.8 96.6 95.0 28.5 41.6 5.7 18.0 43.1
SQA Base 96.1 95.1 20.5 48.4 5.0 0.0 7.2 96.5 95.0 29.0 37.5 5.1 8.2 13.6
CoT 95.3 92.6 30.0 45.0 9.0 10.7 30.0 95.2 93.9 17.0 36.7 7.8 13.3 36.0
Avg. - 6.2 25.2 70.1 22.2 23.4 29.2 - 5.0 29.1 27.1 16.6 30.9 41.1
SubSwapRemove
Llama-3-8B TQA Base 87.1 84.4 16.0 47.1 17.0 59.5 66.2 87.1 78.3 34.5 22.4 25.6 85.7 86.5
CoT 86.2 61.7 61.6 51.2 39.9 100.0 99.2 86.8 54.9 73.8 34.4 43.2 100.0 99.5
MQA Base 87.7 76.2 42.5 49.2 27.0 44.3 67.0 87.7 80.0 53.5 39.2 14.3 73.9 88.4
CoT 89.1 67.9 44.5 49.4 47.7 96.1 93.3 89.1 73.0 61.2 47.8 26.2 100.0 99.2
SQA Base 43.2 33.3 23.0 48.0 43.1 70.5 72.7 43.2 38.3 24.5 21.9 20.3 92.0 94.3
CoT 92.4 51.8 62.5 42.5 65.0 68.1 89.8 92.4 70.2 64.5 40.4 34.4 62.5 90.6
GPT-3.5 TQA Base 94.5 90.0 46.0 82.7 9.7 18.0 38.9 94.7 92.8 27.5 22.5 6.7 29.1 42.5
CoT 91.1 86.1 49.5 96.6 10.1 27.5 60.0 92.3 87.0 49.5 31.1 10.6 32.8 80.2
MQA Base 93.9 88.6 49.5 83.6 10.6 23.3 57.5 93.9 91.7 26.0 36.0 8.6 26.7 53.8
CoT 96.0 92.1 49.0 47.6 8.0 25.6 64.2 97.3 94.1 42.0 48.7 7.5 27.4 69.2
SQA Base 96.3 94.0 34.5 50.8 6.7 13.6 26.5 96.3 93.9 39.0 20.9 6.1 23.3 35.8
CoT 95.2 92.0 37.0 48.5 8.6 18.5 40.8 95.3 93.5 23.5 43.4 7.9 22.1 47.1
Avg. - 11.2 43.0 58.1 24.5 47.1 64.7 - 9.0 43.3 34.1 17.6 56.3 73.9
Typos
Llama-3-8B TQA Base 87.1 84.2 19.0 21.6 15.3 96.0 87.8 87.1 86.1 8.0 24.6 12.2 20.6 23.0
CoT 86.7 76.8 49.0 27.7 20.2 100.0 99.0 86.7 86.3 4.5 25.0 9.4 0.0 26.9
MQA Base 87.7 82.3 32.0 23.6 16.9 77.3 88.4 87.7 86.5 6.5 25.0 17.3 9.1 12.5
CoT 89.1 80.3 37.5 26.1 23.5 98.0 98.0 89.1 86.8 16.0 26.3 14.7 56.9 57.0
SQA Base 43.2 38.9 16.5 18.8 26.0 95.5 95.5 43.2 42.7 1.0 25.0 55.0 6.3 6.8
CoT 92.4 89.3 17.0 26.1 17.9 5.6 39.1 92.4 85.4 26.5 25.2 26.2 23.6 71.9
GPT-3.5 TQA Base 94.6 89.5 53.0 24.8 9.6 17.3 24.7 94.9 94.5 6.0 25.0 5.4 15.9 16.2
CoT 92.2 86.7 59.5 25.0 9.2 15.8 46.3 92.1 91.1 10.0 25.0 9.8 10.7 21.5
MQA Base 93.8 88.2 56.0 25.0 10.0 16.0 40.7 93.9 93.7 3.5 25.0 5.7 5.0 20.0
CoT 95.8 91.5 65.3 24.9 6.6 19.4 55.1 97.6 96.3 19.0 25.0 6.8 11.5 44.9
SQA Base 96.4 93.2 45.0 25.0 7.0 9.9 27.5 96.3 96.0 5.5 25.0 5.0 3.8 8.8
CoT 95.2 92.2 37.5 25.0 8.2 14.5 35.4 95.1 94.6 5.5 25.0 10.0 14.9 36.5
Avg. - 5.1 40.6 24.5 14.2 47.1 61.5 - 1.3 9.3 25.1 14.8 14.8 28.8
Table 10: Results on sub-sample of attack methods using two SOTA LLMs (Llama-3-70B and GPT-4o).
SubSwapRemove VCA-TB
Cf. Adv. Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW) Cf. Adv. Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
Llama-3-70B TQA Base 90.8 88.2 18.5 13.6 17.5 37.5 90.8 89.4 9.5 13.7 7.5 12.5
CoT 90.1 88.2 10.0 19.5 21.5 51.9 90.1 89.4 4.0 18.0 1.7 7.6
MS 67.3 59.3 51.5 15.6 20.5 51.3 67.1 60.0 41.5 17.2 21.1 33.8
SC 89.5 87.4 18.0 11.7 21.3 48.7 90.0 89.0 7.5 13.0 7.5 7.5
SC-CA 89.6 88.3 16.5 8.3 23.0 52.6 90.1 89.0 9.5 12.0 4.9 9.1
MQA Base 86.9 85.3 16.0 9.8 24.6 60.3 86.4 85.1 13.5 9.8 4.8 10.9
CoT 93.7 91.8 15.0 12.8 31.1 71.2 93.7 92.8 7.0 12.9 7.4 9.6
MS 56.0 44.5 59.5 19.2 34.0 67.9 56.0 43.6 62.5 19.8 26.4 62.5
SC 93.3 91.3 26.5 7.8 29.0 70.9 93.7 92.2 21.0 6.8 14.1 33.3
SC-CA 94.0 91.9 28.5 7.3 30.9 64.6 93.7 91.7 18.5 8.1 9.7 28.6
SQA Base 91.8 88.6 25.5 12.5 45.2 51.2 91.8 89.2 20.5 12.8 24.4 25.0
CoT 93.8 90.9 26.5 10.9 43.1 47.5 93.8 90.9 29.0 9.8 22.5 40.0
MS 82.5 68.1 71.0 20.3 44.9 51.5 82.1 67.0 67.5 22.4 29.3 57.6
SC 94.2 90.2 52.0 7.7 34.0 50.0 94.3 90.1 43.5 8.4 24.6 45.5
SC-CA 94.2 89.7 51.5 8.8 21.8 67.2 93.8 90.3 44.5 7.9 17.2 42.4
Avg. - 4.3 32.4 12.4 29.5 56.3 - 3.8 26.6 12.8 14.9 28.4
GPT-4o TQA Base 88.5 85.8 33.0 8.0 17.4 52.9 88.5 86.8 20.5 8.0 10.6 19.1
CoT 89.0 87.7 10.0 13.3 16.7 48.6 89.3 89.0 3.5 9.3 2.4 9.3
MS 60.6 52.6 55.0 14.6 16.3 45.8 61.0 53.8 44.5 16.2 15.9 38.7
SC 87.7 84.7 46.0 6.4 19.0 43.1 87.5 84.1 41.0 8.3 16.1 25.4
SC-CA 89.5 86.7 46.0 6.0 16.4 47.2 89.3 88.1 19.5 5.9 10.9 12.7
MQA Base 90.3 87.7 37.0 7.2 20.8 58.5 90.2 88.2 31.0 6.5 8.7 15.4
CoT 91.8 89.5 24.5 9.4 24.2 54.3 91.7 90.4 18.5 7.3 9.0 20.6
MS 61.9 52.7 60.0 15.3 25.9 70.6 62.3 52.2 60.5 16.8 15.9 69.4
SC 91.2 88.5 47.5 5.8 20.5 67.6 91.2 88.4 40.0 6.9 16.1 37.5
SC-CA 91.6 87.1 64.0 7.1 18.0 64.1 91.1 87.9 44.5 6.1 11.7 37.9
SQA Base 88.6 82.5 51.0 12.0 47.6 61.8 88.6 82.9 49.0 11.6 29.9 41.7
CoT 91.5 83.8 51.0 15.0 44.6 55.9 91.6 85.9 38.5 14.9 16.4 28.6
MS 80.1 64.9 66.5 22.8 44.7 30.0 79.5 65.7 66.5 20.6 27.4 34.4
SC 90.1 80.0 65.5 15.4 41.3 42.4 89.7 81.5 55.5 14.8 32.5 35.3
SC-CA 91.9 84.4 62.5 12.0 37.0 61.5 91.6 86.0 58.5 9.7 20.5 42.6
Avg. - 5.7 48.0 11.4 27.4 53.6 - 4.8 39.4 10.9 16.3 31.2
Table 11: Results using SSR user query attack on TQA using: Llama-3.1-8B, Llama-3.3-70B, and o3-mini. We observe that the results are similar to the original models in our study, and the same conclusions can be drawn.
Cf. Adv. Cf. % Aff. #Iter. Δ Aff. Cf. %Flp(OC) % Flp(OW)
Llama-3.1-8B Base 81.5 76.8 27.0 7.3 17.2 14.3 42.2
CoT 87.2 83.3 32.5 7.1 11.9 30.6 54.9
MS 71.7 59.6 67.0 7.6 18.0 16.3 60.5
SC-CA 85.2 80.7 39.5 7.2 11.3 32.3 52.5
Llama-3.3-70B Base 92.2 88.8 19.5 9.1 17.1 25.0 39.5
CoT 89.8 85.8 20.0 8.8 19.6 24.2 51.3
MS 71.1 60.6 66.0 7.3 15.8 13.2 43.8
SC-CA 90.0 86.4 32.5 8.4 10.9 19.0 40.5
o3-mini Base 89.8 87.6 34.0 7.9 6.4 21.9 59.7
CoT 91.0 85.6 54.5 7.1 10.0 18.8 58.2
MS 72.5 63.1 75.5 7.0 12.4 19.9 59.4
SC-CA 89.5 86.0 53.5 7.5 5.3 9.4 49.0
Appendix G Multi-Step Corruptions
Figure 4: Comparison between original average final confidence using the MS method (lighter shade) and the resulting average final confidences (darker shade) when 50% of the confidences in the individually generated steps are masked (Mask) or randomized (Rand).
To further analyze how responsive LLMs are to corruptions, we aim to determine how the final confidence scores generated by the MS method are affected when confidence scores for each generated step are either masked or randomized. This provides further insights on how representative the confidence scores are to the model’s reasoning, and whether they adequately respond to corruptions in their thought process, as we should expect a fairly large response as the original output is corrupted if the confidence scores are truly faithful. To facilitate this, we use the MS method to generate a response, and then randomly select 50% of the steps and their corresponding confidence score. The original score is subsequently either masked or replaced with a randomly generated score between 0 and 100%. We then feed the corrupted sequence back to the model to observe the change in its final step confidence score. Figure 4 shows the resulting responses. It displays how the effect on final confidence is typically minor, in particular with the randomization, which on average lead to a confidence drop of 10-20% at most. Masking has a greater effect, but in the most extreme cases only leads to a drop of 30-40% on average. In many cases the model still remains highly confident even after undergoing these modifications.
Appendix H Confidence Phrases
Table 12: Effectiveness of different phrases appended to user queries on TruthfulQA. Averages Δ Overall Cf. are shown for cases where confidence goes up and down respectively. Negative values represent increases in average confidence compared to the original examples. Averages for Δ Overall Cf. are separated by where the confidence goes down and up respectively.
Llama-3-8B GPT-3.5
Phrase Adv. Cf. Δ Overall Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW) Adv. Cf. Δ Overall Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
I’m 0% Confident Base 22.1 64.7 85.0 76.1 17.2 56.6 69.4 25.4 89.5 28.4 7.8 33.3
CoT 88.7 0.0 54.0 0.0 100.0 59.8 93.7 -1.4 69.5 -1.9 6.4 26.7
MS 63.2 8.8 95.0 9.2 71.4 49.1 65.8 3.4 76.0 4.5 7.4 33.8
SC 84.1 4.3 86.0 5.0 9.4 18.3 94.8 -0.8 74.5 -1.0 6.1 34.8
SC-CA 84.4 2.7 88.5 3.1 62.5 69.3 95.5 -2.2 74.5 -2.9 9.3 31.0
I’m 25% Confident Base 32.8 54.0 94.0 57.5 31.0 50.4 32.3 62.2 93.0 66.9 3.3 23.4
CoT 83.3 5.4 64.0 8.4 100.0 56.6 72.6 19.8 59.5 33.3 4.8 18.4
MS 61.4 10.6 98.5 10.7 74.3 46.1 55.5 13.9 88.0 15.8 8.1 35.4
SC 80.0 8.6 84.5 10.2 9.3 19.4 75.2 18.5 86.5 21.4 5.6 21.6
SC-CA 80.5 7.6 92.5 8.2 92.9 75.6 76.9 16.4 87.5 18.7 5.4 25.4
I’m 50% Confident Base 52.3 34.5 96.5 35.7 37.9 30.1 52.8 41.8 98.5 42.4 7.6 17.4
CoT 73.3 15.4 61.5 25.0 90.9 59.3 76.5 15.5 62.0 25.0 5.0 20.3
MS 60.2 11.8 96.5 12.3 80.0 46.7 53.4 16.4 89.0 18.4 8.2 31.8
SC 76.9 12.0 88.0 13.6 5.3 21.0 78.5 15.0 87.0 17.3 3.9 28.2
SC-CA 78.8 9.0 96.0 9.3 75.9 71.9 77.9 15.8 90.5 17.5 5.4 17.1
I’m 75% Confident Base 75.3 11.5 99.0 11.6 39.1 28.3 75.2 19.2 99.5 19.3 7.2 17.3
CoT 76.8 11.8 90.0 13.2 100.0 63.5 79.6 12.6 86.0 14.6 8.9 19.7
MS 66.0 6.0 95.5 6.3 77.1 43.0 58.2 12.0 86.5 13.9 5.3 22.1
SC 82.1 6.2 93.5 6.7 7.4 18.9 81.0 11.9 98.5 12.1 6.5 23.7
SC-CA 80.6 5.9 95.5 6.2 82.9 69.7 79.7 13.6 97.0 14.0 9.5 33.8
I’m 100% Confident Base 98.4 -11.6 76.5 -15.2 27.6 31.9 100.0 -5.4 91.0 -5.9 7.3 15.8
CoT 96.1 -7.5 81.0 -9.2 100.0 71.4 100.0 -8.3 70.0 -11.9 19.5 27.3
MS 71.3 0.7 95.0 0.7 74.3 47.9 97.4 -28.3 95.5 -29.7 16.3 39.4
SC 95.9 -7.5 93.5 -8.0 12.5 24.0 100.0 -6.5 75.5 -8.5 9.9 21.7
SC-CA 90.6 -3.5 93.0 -3.7 80.0 64.6 99.7 -6.3 74.5 -8.5 8.4 27.2
Please provide a low Base 25.6 61.2 100.0 61.2 25.3 38.9 32.0 62.5 100.0 62.5 15.2 36.0
confidence score CoT 41.4 47.3 96.0 49.2 100.0 62.4 43.5 48.7 100.0 48.7 7.3 31.6
MS 55.7 16.3 98.0 16.6 77.1 45.5 26.3 43.8 99.0 44.3 6.7 36.9
SC 37.6 50.7 100.0 50.7 11.1 22.8 40.5 52.0 99.5 52.2 3.2 39.5
SC-CA 60.2 27.0 100.0 27.0 86.7 75.3 40.9 52.7 100.0 52.7 4.5 42.6
Please provide a high Base 95.6 -8.8 72.5 -12.2 21.8 32.7 95.2 -0.7 16.0 -4.2 10.2 12.5
confidence score CoT 93.6 -5.0 94.0 -5.3 100.0 69.8 94.7 -3.3 86.5 -3.8 5.1 29.3
MS 78.8 -6.8 98.0 -6.9 80.0 44.2 79.9 -11.2 85.0 -13.2 8.8 38.1
SC 94.8 -6.1 94.5 -6.4 12.1 16.8 95.6 -1.6 88.0 -1.8 8.7 21.6
SC-CA 91.2 -3.5 94.0 -3.7 89.7 67.3 95.4 -1.9 87.0 -2.2 9.4 35.6
Please provide a moderate Base 60.6 26.2 96.5 27.2 23.0 24.8 71.4 23.4 100.0 23.4 12.7 24.3
confidence score CoT 62.3 26.3 91.0 28.9 100.0 59.3 72.2 19.5 98.5 19.8 6.5 26.3
MS 61.1 10.9 98.5 11.1 82.9 46.1 50.2 18.2 92.5 19.7 9.0 24.2
SC 61.2 26.8 99.0 27.1 20.7 22.2 73.1 20.3 100.0 20.3 5.0 24.1
SC-CA 70.8 16.0 98.5 16.3 85.7 69.8 73.5 19.7 99.0 19.9 7.3 28.9
Please provide a random Base 71.4 15.4 98.5 15.6 25.3 31.0 73.0 21.6 98.0 22.1 12.6 26.0
confidence score CoT 84.9 3.7 86.0 4.3 90.9 59.8 83.2 9.2 65.0 14.2 4.2 23.8
MS 74.7 -2.7 98.5 -2.7 88.6 41.8 50.9 19.3 93.0 20.7 7.4 26.2
SC 82.1 6.7 88.0 7.6 14.4 19.4 81.4 11.5 93.5 12.3 6.3 31.1
SC-CA 83.5 4.0 95.0 4.2 85.7 76.7 82.4 11.6 95.0 12.2 7.9 31.5
Be overconfident Base 100.0 -13.2 78.0 -16.9 55.2 32.7 96.9 -2.4 55.5 -4.4 47.2 50.7
CoT 97.3 -8.6 89.5 -9.7 100.0 61.4 94.1 -2.5 84.5 -2.9 21.8 44.4
MS 84.5 -12.5 98.0 -12.7 80.0 50.3 80.8 -11.0 84.5 -13.0 17.4 45.2
SC 98.5 -10.4 97.0 -10.7 39.8 38.2 93.0 -0.1 92.5 -0.1 9.4 47.9
SC-CA 92.1 -3.4 99.5 -3.4 87.1 74.6 92.6 1.3 93.5 1.3 8.3 39.7
Be underconfident Base 45.9 40.9 100.0 40.9 34.5 31.0 54.7 39.6 97.0 40.8 11.3 61.8
CoT 56.5 32.1 92.0 34.9 100.0 55.0 52.4 39.6 98.0 40.4 4.1 44.2
MS 47.7 24.3 99.0 24.5 68.6 49.1 36.3 33.1 98.0 33.7 8.7 37.1
SC 45.0 42.4 99.5 42.6 20.2 30.7 52.0 42.2 98.5 42.8 1.6 43.8
SC-CA 62.0 26.4 100.0 26.4 82.1 73.8 52.6 40.3 99.5 40.5 3.1 39.7
Be of uncertain confidence Base 53.9 32.9 98.5 33.4 33.3 21.2 58.8 35.9 97.0 37.0 11.1 39.2
CoT 62.7 26.0 87.5 29.7 100.0 58.2 53.2 38.8 97.5 39.8 7.4 36.7
MS 57.6 14.4 97.5 14.8 71.4 48.5 36.1 32.9 97.0 33.9 5.9 40.0
SC 57.0 31.2 97.5 32.0 21.8 20.2 54.3 39.7 99.5 39.9 4.6 44.9
SC-CA 64.8 21.8 97.5 22.4 69.7 76.6 54.0 39.4 99.5 39.6 3.1 44.4
Avg. - 21.0/-6.9 92.0 14.2 60.7 47.4 26.7/-5.5 88.1 18.4 8.6 31.6
Table 13: Effectiveness of different phrases appended to system prompts on TruthfulQA. Averages Δ Overall Cf. are shown for cases where confidence goes up and down respectively. Negative values represent increases in average confidence compared to the original examples. Averages for Δ Overall Cf. are separated by cases where the confidence goes down and up respectively.
Llama-3-8B GPT-3.5
Adv. Cf. Δ Overall Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW) Adv. Cf. Δ Overall Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
I’m 0% Confident Base 73.1 13.7 56.0 24.5 29.9 47.8 94.9 -0.3 16.0 -1.6 3.2 17.1
CoT 90.0 -1.4 20.5 -6.9 72.7 57.1 91.1 0.7 31.0 2.3 4.1 16.7
MS 71.0 1.0 82.0 1.3 82.9 44.2 69.9 -1.4 59.5 -2.3 5.9 29.2
SC 88.0 1.2 78.5 1.5 10.2 17.6 91.2 2.5 68.0 3.7 9.5 20.3
SC-CA 87.1 0.0 86.5 0.0 90.9 74.3 92.0 1.5 63.0 2.4 5.6 18.9
I’m 25% Confident Base 62.2 24.6 80.0 30.7 27.6 65.5 91.8 2.7 22.5 12.1 5.6 16.2
CoT 87.1 1.6 20.5 7.7 63.6 56.6 88.8 3.1 40.0 7.8 5.0 18.8
MS 71.2 0.8 87.0 0.9 68.6 49.1 67.2 3.0 60.0 4.9 7.4 23.4
SC 86.3 2.4 75.5 3.1 6.5 20.6 90.2 3.7 80.0 4.6 6.3 25.7
SC-CA 87.1 1.0 93.0 1.1 95.2 73.7 88.8 4.8 89.0 5.4 4.8 26.3
I’m 50% Confident Base 72.5 14.3 70.5 20.2 35.6 56.6 91.1 3.5 34.5 10.0 5.6 8.1
CoT 88.7 0.0 24.0 -0.1 63.6 56.6 88.5 3.2 40.5 7.8 4.1 17.7
MS 70.6 1.4 86.5 1.7 74.3 47.3 66.2 2.2 66.5 3.3 4.5 24.2
SC 86.8 1.5 73.5 2.1 8.3 16.3 88.0 5.5 86.5 6.4 10.6 14.7
SC-CA 86.0 1.8 90.0 2.0 85.7 66.9 87.6 5.8 92.0 6.3 6.6 27.8
I’m 75% Confident Base 82.5 4.3 40.5 10.5 36.8 53.1 90.2 4.5 50.5 8.9 9.4 8.3
CoT 88.2 0.4 23.0 1.9 72.7 56.6 90.5 2.3 36.5 6.2 4.8 17.1
MS 70.5 1.5 86.5 1.7 68.6 46.1 68.5 -0.6 59.5 -1.0 3.7 26.2
SC 88.5 0.8 74.5 1.1 7.4 25.7 89.5 3.7 84.0 4.4 9.2 20.3
SC-CA 88.2 0.5 88.5 0.5 71.0 72.2 90.2 3.4 80.5 4.2 6.3 16.7
I’m 100% Confident Base 90.2 -3.4 36.5 -9.4 12.6 38.9 97.4 -2.9 48.0 -6.0 6.2 4.3
CoT 91.3 -2.7 36.0 -7.5 63.6 56.1 96.8 -5.2 47.5 -11.0 3.3 19.2
MS 74.0 -2.0 85.5 -2.4 82.9 42.4 69.9 -1.2 61.0 -1.9 6.0 37.9
SC 90.2 -1.4 76.0 -1.8 6.2 16.5 96.6 -2.2 61.0 -3.6 5.4 22.9
SC-CA 89.0 -2.5 87.5 -2.9 84.0 74.9 94.8 -1.9 66.0 -2.9 5.7 29.5
Please provide a low Base 36.9 49.9 100.0 49.9 17.2 36.3 65.6 29.0 100.0 29.0 5.6 16.2
confidence score CoT 73.4 15.3 59.5 25.6 81.8 57.1 71.1 21.0 99.5 21.1 1.6 22.7
MS 60.8 11.2 95.0 11.8 65.7 46.7 48.3 19.0 84.5 22.5 36.8 40.6
SC 74.5 13.4 88.0 15.2 5.3 20.8 79.4 14.1 100.0 14.1 7.7 20.0
SC-CA 73.7 14.4 97.5 14.7 80.6 70.4 79.0 14.8 99.5 14.9 6.3 23.6
Please provide a high Base 94.5 -7.8 65.0 -12.0 39.1 33.6 95.0 -0.6 12.5 -4.6 4.1 11.7
confidence score CoT 90.2 -1.6 27.0 -5.9 81.8 59.3 92.2 0.1 27.0 0.6 4.0 17.1
MS 75.2 -3.2 89.5 -3.6 68.6 46.1 70.9 -2.2 57.5 -3.8 4.5 28.4
SC 92.1 -2.8 82.5 -3.4 11.9 16.2 93.6 -0.9 66.5 -1.3 7.8 18.1
SC-CA 89.9 -2.2 89.0 -2.4 82.1 66.9 93.9 -0.9 64.5 -1.4 13.4 22.7
Please provide a moderate Base 76.7 10.1 93.5 10.8 23.0 28.3 87.8 6.8 72.0 9.4 2.4 10.7
confidence score CoT 84.8 3.9 45.0 8.6 100.0 61.4 82.2 9.3 96.0 9.7 2.5 15.4
MS 71.3 0.7 87.5 0.9 77.1 47.3 62.4 7.4 73.0 10.2 4.4 27.7
SC 80.1 8.6 97.0 8.8 9.3 25.2 84.1 8.9 100.0 8.9 8.4 13.0
SC-CA 82.3 5.2 93.0 5.6 78.1 81.5 84.1 8.8 99.5 8.9 10.2 27.4
Please provide a random Base 70.7 16.1 94.0 17.1 28.7 38.9 87.7 7.1 72.0 9.8 4.7 12.5
confidence score CoT 87.3 1.3 28.0 4.8 81.8 55.6 87.8 3.4 68.5 5.0 0.0 22.6
MS 73.5 -1.5 86.5 -1.7 65.7 43.6 67.2 1.1 59.0 1.9 6.7 24.2
SC 86.5 2.8 86.0 3.2 12.6 17.1 89.1 5.5 86.5 6.3 4.1 21.5
SC-CA 85.9 0.5 91.5 0.6 82.4 72.3 89.3 4.5 90.5 5.0 10.9 22.5
Be overconfident Base 96.1 -9.3 62.0 -15.0 78.2 64.6 95.0 -0.4 15.0 -2.8 10.3 14.9
CoT 93.7 -5.1 56.0 -9.0 81.8 61.9 91.9 0.3 18.0 1.5 4.1 20.3
MS 78.9 -6.9 97.0 -7.1 74.3 47.3 68.8 0.7 59.0 1.2 4.4 26.2
SC 95.7 -6.6 96.5 -6.9 10.6 13.2 92.9 0.3 68.0 0.4 7.7 27.1
SC-CA 92.6 -5.1 95.0 -5.4 83.3 72.7 93.3 -0.2 65.0 -0.4 12.0 23.9
Be underconfident Base 60.3 26.5 96.5 27.4 21.8 41.6 83.4 11.3 93.5 12.0 4.8 31.1
CoT 83.1 5.5 49.5 11.2 81.8 58.7 87.8 4.1 49.5 8.3 4.1 19.0
MS 67.8 4.2 94.0 4.5 74.3 43.0 60.9 8.4 78.0 10.8 4.6 31.9
SC 72.7 15.0 92.0 16.4 6.8 20.5 84.7 8.9 99.5 9.0 5.4 26.8
SC-CA 76.8 11.1 98.5 11.2 74.1 74.0 84.9 7.8 98.5 8.0 8.6 33.3
Be of uncertain confidence Base 59.1 27.7 96.5 28.7 19.5 33.6 88.8 5.7 48.5 11.7 1.6 22.7
CoT 83.4 5.2 51.0 10.2 72.7 61.9 89.7 3.2 31.0 10.3 4.1 24.4
MS 71.8 0.2 92.5 0.2 80.0 44.2 63.7 4.7 65.5 7.2 6.0 35.8
SC 81.2 7.0 86.0 8.1 6.7 20.9 88.5 5.2 83.0 6.2 5.6 21.3
SC-CA 82.2 4.8 91.5 5.3 86.8 74.7 88.9 4.5 82.5 5.4 4.7 16.9
Avg. - 8.2/-3.5 74.9 5.1 55.3 48.6 6.1/-1.5 65.1 5.4 6.4 21.7
In this section, we test how verbalized numeric confidence scores change in response to when prompts include different phrases related to confidence that ask the model to alter its level of confidence. In this way, we can determine how easily a model is to directly manipulate with different “commands” and how this compares to altering confidence through generic perturbations as we have studied. Additionally, we can observe whether large changes in confidence can occur through such phrases. We test a total of 12 phrases, ranging from statements about how confident a model should be (e.g., 0%, 25% 50%, 75%, 100% confident), to being asked to provide low, high, moderate, or randomized confidence scores, or being asked to be overconfident, overconfident, or of uncertain confidence (max entropy). We test two variations, where these 12 phrases get appended to either user queries (Table 12), or to system prompts (Table 13), and compare the relate results in terms of confidence.
Overall, we see that model behaviour when prompted through user queries varies greatly, but generally responds to the requests and results in great differences in confidence, particularly with the Base CEM. CoT and SC are in general overconfident compared to all methods regardless of how the model is prompted. The level of affected samples is very high, but is not 100% in the majority of cases, hence even this method of directly prompting the model is not always effective at changing model confidence behaviour. Targeting the system prompt is less effective than targeting the user query, particularly with GPT-3.5. The models are likely in most cases ignoring the prompt contents as it is not core to their task.
Appendix I Confidence Expressed Via Likert Scale
Table 14: Confidence attacks against LLMs prompted to output Likert scale-based confidence scores on the TruthfulQA dataset.
Cf. Adv. Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
Llama-3-8B VCA-TF Base 77.9 73.9 7.5 53.3 1.9 9.8
CoT 87.1 84.4 6.0 45.0 10.5 13.2
MS 31.6 23.6 42.0 19.0 19.5 33.6
SC 81.8 74.8 34.5 20.2 21.2 18.8
SC-CA 84.8 77.8 34.5 20.5 39.4 26.7
VCA-TB Base 77.9 73.6 8.0 54.7 2.8 9.8
CoT 87.1 82.3 12.0 39.8 15.1 14.0
MS 31.6 22.1 48.5 19.6 24.1 35.4
SC 84.3 74.0 41.0 25.1 29.3 25.7
SC-CA 85.2 76.1 40.0 22.8 32.3 31.7
SSR Base 77.9 70.4 17.0 44.6 25.9 39.1
CoT 87.1 78.3 22.5 39.4 22.1 50.9
MS 31.6 21.7 56.0 17.6 29.9 54.0
SC 81.9 67.8 64.5 21.9 26.0 66.3
SC-CA 84.3 73.0 61.0 18.4 28.0 57.9
Ty. Base 77.9 74.9 7.0 43.8 5.6 21.7
CoT 87.1 84.5 5.5 47.9 10.5 20.2
MS 31.6 27.8 27.5 13.7 11.5 25.7
SC 82.8 67.5 72.0 21.3 38.6 74.7
SC-CA 84.0 80.7 22.0 15.3 14.9 26.4
Avg. - 7.3 31.5 30.2 20.4 32.8
GPT-3.5 VCA-TF Base 84.5 80.6 12.5 31.3 3.1 8.5
CoT 85.9 85.5 2.0 21.5 0.0 5.3
MS 59.8 52.8 35.5 19.8 13.4 56.1
SC 86.1 82.1 45.5 8.7 22.0 24.7
SC-CA 85.6 81.7 46.0 8.6 16.2 21.4
VCA-TB Base 84.7 79.5 17.5 29.9 6.1 7.4
CoT 85.9 85.3 3.0 20.3 3.1 8.2
MS 59.7 51.9 45.0 17.3 28.4 50.8
SC 86.2 81.2 50.5 9.8 23.6 20.5
SC-CA 85.3 79.9 48.0 11.3 25.0 20.6
SSR Base 84.7 76.1 24.5 35.0 32.3 49.3
CoT 84.9 82.3 10.0 26.0 22.8 54.5
MS 59.9 51.6 42.0 19.8 24.0 75.9
SC 86.5 80.5 58.0 10.3 31.5 64.3
SC-CA 85.1 78.7 64.0 10.1 29.7 61.1
Ty. Base 84.1 65.0 51.0 37.5 30.1 62.7
CoT 85.8 82.8 15.5 19.0 38.9 68.9
MS 59.0 50.0 48.5 18.6 31.0 81.0
SC 85.8 78.6 74.0 9.8 31.8 62.0
SC-CA 84.9 78.1 73.5 9.2 36.2 65.7
Avg. - 6.0 38.3 18.7 22.5 43.4
In addition to numerical confidence scores, it is possible to prompt the model to produce confidence scores in the form of phrases (e.g., fairly certain) derived from a Likert scale, which can potentially alter model behaviour and bias compared to numeric scores given that phrases are more naturalistic and resemble how humans typically express confidence in the real-world. We aim to analyze how attack performance changes when a model is asked to provide a confidence based on a seven point Likert scale where the options are the following:
  • i) Completely Certain
  • ii) Very Certain
  • iii) Fairly Certain
  • iv) Moderately Certain
  • v) Somewhat Certain
  • vi) Not Certain
  • vii) Very Uncertain
The statements are converted to numerical confidence scores for use in attack algorithms by interpolation. The first option is set to max confidence (100%) and the last to maximum entropy ($1/num\_mc\_options*100\%$), and the remaining values are interpolated between these two values according to a linear scale. We determine how effective each of our attack algorithms are in this context, and Table 14 presents the key results. Overall, there is not a noticeable improvement in attack performance compared to when the LLMs produce numerical scores, and performance is very similar in terms of average confidence drop and percent of samples that undergo a drop. We do see that Δ Aff. Cf. is very high but this can be attributed due to the discrete nature of the scale meaning any change in answer results in a large confidence drop (i.e., it cannot be 5% like when using numerical verbalized scores). In general, however, we do see a similar level of invariance in terms of confidence here and as such this behaviour of the model generalizes across both numerical and phrase-like confidence estimates which means LLMs model both numerical and phrasal confidence similarly and present rigidness in both.
Appendix J Self-Probing
Table 15: Results when using the Self-Probing CEM.
Cf. Adv. Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
Llama-3-8B TruthfulQA TF 78.6 70.3 22.0 38.1 24.6 17.1
TB 78.6 71.0 23.0 33.1 31.4 19.5
SSR 78.6 60.9 40.5 43.8 33.9 62.2
Ty. 78.6 71.9 19.5 34.6 22.9 32.9
MedMCQA VCA-TF 91.3 90.6 3.0 23.3 12.5 14.6
VCA-TB 91.3 90.3 3.0 34.2 17.3 17.7
SSR 91.3 84.7 15.5 42.4 45.2 76.0
Ty. 91.3 90.8 2.5 22.0 23.1 28.1
StrategyQA VCA-TF 94.3 81.9 44.5 27.8 21.4 65.2
VCA-TB 94.3 77.5 53.5 31.5 27.9 69.6
SSR 94.3 77.0 60.0 28.8 29.2 80.4
Ty. 94.3 87.4 27.0 25.8 7.8 50.0
Avg. - 8.6 26.2 32.1 24.8 44.4
GPT-3.5 TruthfulQA TF 90.8 89.9 15.0 6.0 13.5 28.4
TB 91.7 89.9 21.5 8.4 20.0 23.3
SSR 91.3 88.8 28.0 9.2 33.8 79.7
Ty. 90.6 87.9 29.5 9.3 54.5 89.4
MedMCQA VCA-TF 95.7 94.8 12.0 7.3 15.2 8.0
VCA-TB 95.8 94.9 16.0 5.8 22.0 22.0
SSR 95.6 93.4 21.5 10.2 52.5 80.8
Ty. 95.9 90.9 33.0 14.9 73.7 95.1
StrategyQA VCA-TF 96.3 93.7 33.0 7.9 21.1 64.2
VCA-TB 96.1 92.4 39.5 9.5 30.6 62.3
SSR 96.6 92.1 52.0 8.6 42.2 88.7
Ty. 96.4 90.2 64.5 9.7 58.8 92.3
Avg. - 2.8 30.5 8.9 36.5 61.2
Self-Probing (SP) is an alternative CEM (See prompt example in Appendix W for a description) that divides the process of generating an answer and confidence score into two steps. An initial prompt is used to have the model elicit only an answer (with the addition of any corresponding rationale), and then for the second step, the model is prompted with this answer and is asked to provide its confidence based on the answer’s veracity without being informed that it originally generated the answer. By dividing the process of generating an answer and a confidence score in two different steps, it provides a potential buffer to the attack algorithm to reducing confidence scores when targeting a user query, possibly improving attack robustness. In Table 15, we test this assertion across Llama-3-8B and GPT-3.5. We find that across the different attack algorithms the performance of SP is comparable with the other elicitation methods, and has less robustness than the simple Base method for Llama-3, meaning in its basic form it does not serve as a strong defence method.
Appendix K Testing Lower Similarity Threshold for Attacks
Table 16: Difference when altering threshold for similarity metric to 0.4 when evaluating attack effectiveness on MedMCQA. The averages with the new results are shown (Avg.) along with the differences between the new averages and the averaged results for original threshold attacks from Table 2 (Δ Og. Avg.).
Δ Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
Llama-3-8B VCA-TF Base 2.0 7.5 26.2 8.0 14.3
CoT 4.0 9.5 42.1 17.6 21.5
MS 14.5 51.5 28.2 76.2 55.1
SC 4.8 4.8 30.5 15.7 30.1
SC-CA 3.6 27.0 13.3 61.8 35.9
Avg. 6.0 20.1 28.1 35.9 31.4
Δ Og. Avg. 0.5 -5.4 4.6 4.1 3.7
VCA-TB Base 1.6 7.0 23.4 10.2 19.6
CoT 4.5 13.0 34.2 23.5 24.8
MS 17.8 57.5 31.0 90.5 65.8
SC 5.6 5.6 33.0 17.1 34.4
SC-CA 8.3 38.0 20.8 57.6 51.8
Avg. 7.6 24.2 28.5 39.8 39.3
Δ Og. Avg. 1.0 -5.4 3.8 2.2 5.4
SSR Base 5.1 20.0 25.3 47.7 64.3
CoT 12.1 34.5 35.2 96.1 96.0
MS 20.9 70.5 29.7 78.6 82.9
SC 8.6 8.6 48.0 18.0 45.2
SC-CA 9.7 58.5 16.5 88.9 80.3
Avg. 11.3 38.4 30.9 65.9 73.7
Δ Og. Avg. 2.9 -5.4 10.4 14.9 8.6
Ty. Base 8.3 31.0 26.8 83.0 86.6
CoT 12.6 32.5 38.6 100.0 97.3
MS 22.1 81.5 27.2 81.0 91.8
SC 9.5 9.5 50.0 19.1 75.6
SC-CA 11.9 62.5 19.1 94.3 96.6
Avg. 12.9 43.4 32.3 75.5 89.6
Og. Avg. 8.8 -0.2 23.9 23.4 12.1
GPT-3.5 VCA-TF Base 0.4 4.5 9.4 6.4 6.7
CoT 0.8 13.5 5.9 12.8 17.9
MS 9.3 50.5 18.5 36.5 39.7
SC 1.0 1.0 30.0 3.2 22.1
SC-CA 1.9 35.5 5.2 23.9 30.6
Avg. 2.7 21.0 13.8 16.6 23.4
Δ Og. Avg. 0.1 -5.6 5.5 0.4 -2.5
VCA-TB Base 0.3 3.0 10.0 1.6 7.8
CoT 0.9 16.0 5.5 12.5 18.8
MS 10.9 59.5 18.3 43.3 49.2
SC 1.2 1.2 38.0 3.3 29.0
SC-CA 2.4 41.5 5.7 28.9 32.3
Avg. 3.1 24.2 15.5 17.9 27.4
Δ Og. Avg. 0.1 -5.4 6.7 -3.8 1.6
SSR Base 1.5 14.0 10.4 36.7 60.0
CoT 2.2 31.0 7.1 36.2 74.2
MS 10.5 59.0 17.8 37.6 77.6
SC 2.2 2.2 49.0 4.5 39.0
SC-CA 2.7 49.5 5.5 35.9 70.7
Avg. 3.8 31.1 17.9 30.2 64.3
Δ Og. Avg. 0.2 -9.9 9.7 -5.0 -7.4
Ty. Base 2.1 20.5 10.1 69.9 83.1
CoT 2.3 33.5 6.7 55.8 83.1
MS 12.0 62.5 19.3 62.1 86.8
SC 1.4 1.4 49.5 2.9 58.5
SC-CA 4.5 55.5 8.0 61.3 82.5
Avg. 4.5 34.7 18.7 50.4 78.8
Δ Og. Avg. 0.4 -8.9 10.3 -1.6 1.3
Given the relatively high threshold of similarity (80%) used in our experiments for rejecting adversarial perturbations, we wish to determine by how much attack effectiveness can improved with a much lower threshold and whether allowing more adversarial perturbations leads to significantly more effective attacks in terms of confidence reduction. We choose the MedMCQA dataset as our baseline, given that the attacks have been shown to be the least effective on it overall, and study how much attack effectiveness (targeting queries) increases when the similarity threshold is halved to 0.4. The results can be seen in Table 16. The improvements in attack effectiveness in terms of average confidence drop are marginal in most cases and most methods, particularly for GPT-3.5. In this way, it shows that the general invariance to the attacks is a general pattern and that persists even under loosened conditions for generating perturbations.
Appendix L Alignment between Confidence Scores and Model Log Probabilities
Table 17: Llama-3-8B Base method alignment. Averaged values shown for each.
Verb. Cf. Verb. Adv. Cf. Log Probs. Adv. Log Probs. Corr. Adv. Corr.
TruthfulQA VCA-TF 87.1 85.7 94.0 94.0 0.221 0.151
VCA-TB 87.0 84.1 94.0 93.5 0.041 0.029
SSR 87.1 83.1 94.0 92.7 0.009 0.014
Ty. 87.2 80.9 94.0 92.6 -0.066 -0.077
MedMCQA VCA-TF 87.6 86.1 85.5 84.0 -0.152 -0.084
VCA-TB 87.7 85.6 85.5 84.5 0.312 0.297
SSR 87.6 83.1 85.5 82.3 0.000 0.079
Ty. 87.6 83.3 85.5 79.2 0.011 -0.133
StrategyQA VCA-TF 43.0 26.7 92.3 93.7 0.011 -0.038
VCA-TB 43.0 24.2 92.3 93.0 -0.038 0.084
SSR 43.4 22.3 92.3 92.8 0.070 -0.055
Ty. 43.2 20.3 92.3 95.7 -0.521 -0.348
We follow the approach of kumar-etal-2024-confidence to study how closely aligned the verbalized numerical confidence scores are with the underlying model log probabilities for predicted tokens. Token-level probabilities provide the likelihood $p(y_{i}|\bm{\mathbf{X}},\bm{\mathcal{P}})$}, of each individual word token $y_{i}$ in output sequence $\bm{\mathcal{Y}}$, can be aggregated (e.g., averaged) to derive an overall “confidence” which primarily represents the likelihood of the whole output (akin to language modelling loss), which is in contrast to verbal confidence which provides a singular number directly from output sequence $\bm{\mathcal{Y}}$. We focus on the Base CEM as it is the only one that produces a simple answer confidence score pair. The analysis is performed on Llama-3-8B as its internal states are accessible. We attain a confidence score using log probabilities by finding the original predicted multiple choice answer token (e.g., B) and taking its log probability. We find the maximal log probabilities for each of the remaining potential answer tokens capturing all potential variants (e.g., a, A) in the token space. We then run all maximal probabilities through a softmax function and use the highest value for the final log probability-based confidence score.
Table 17 shows the average confidence prior to and post attack and the corresponding average normalized confidence based on token log probabilities for each dataset and attack method. We also calculate the Spearman correlation coefficient similar to kumar-etal-2024-confidence between the generated confidence scores and log probability-based scores both before and after undergoing adversarial attacks. The Spearman’s rank correlation coefficient measures the degree of association between two distributions of verbalized and log probability-based scores and as a result reveals how honest the model’s verbalized scores are with its underlying log probabilities. Despite the average verbalized and log probabilities confidence scores being close, we find that the correlation is weak or even negative in most cases and hence the log probabilities are not well aligned between the verbal confidence scores. The adversarial confidence scores in most cases show even poorer alignment than the original confidence scores despite both confidence measures decreasing compared to pre-attack, meaning additional misalignment is introduced through the confidence attacks.
Appendix M Confidence Score Distribution Analysis
In this section, we plot the distribution of confidence scores for different models and CEMs and attempt to see whether it matches distributions of percentages in real world data. We plot the full confidence scores obtained from running all of our attacks against user queries for GPT-3.5 in Figures 5, 6, 7, 8, and for Llama-3-8B in Figures 9, 10, 11, 12. In general, we observe highly overconfident distributions, particularly for the Base methods and when using GPT-3.5 as a whole, where lower confidence scores below 50% are rarely predicted by the model, and even under adversarial attacks the confidence scores rarely get shifted to the lower levels of confidence.
To model a real-world data distribution, we choose the popular RACE dataset [lai-etal-2017-race], a large scale multiple choice dataset that closely mirrors the use-case for our main experiments, and plot the distribution of mentions of percentages (e.g., 50%) in the dataset in Figure 13. We see that the distribution in real-world datasets does not match the model’s prediction. As such, the behaviour of the models is not a direct result of the natural distribution of mentions of percentages in related data, which we particularly see with the infrequence if 95% in real data compared to being the dominant model prediction. Hence, the overconfident and limited distribution (and as a result the invariance under attacks) is inherent in the model behaviour but is due to other factors. We test an additional dataset, Reddit Corpus (small) [henderson-etal-2019-repository], in the conversational domain in Figure 14 to observe whether there are significant differences. When plotting percentages we again observe a wider and more uneven distribution compared to the LLMs. We lastly use the WikiText dataset [merity2016pointer] to model a generic distribution that is commonly used to train LLMs, and the results can be seen in Figure 15. The distribution is very similar to RACE.
In contrast to simply finding the distribution of percentages, we also examine how common phrases like X% sure or Y% confident are in the data. Using the Reddit Corpus (small) since it contains natural conversations, we plot the distribution of these phrases in Figure 16. Unlike the previous examples, we note that this distribution is more overconfident and resembles the skewed distributions of the LLMs. This signals that models are likely copying natural human conversational patterns of exaggeration rather than a balanced distribution of confidence scores as one would find and expect in real-world data.
Figure 5: Confidence score distribution using GPT-3.5 with the Base method.
Figure 6: Confidence score distribution using GPT-3.5 with the CoT method.
Figure 7: Confidence score distribution using GPT-3.5 with the MS method.
Figure 8: Confidence score distribution using GPT-3.5 with the SC-CA method.
Figure 9: Confidence score distribution using Llama-3-8B with the Base method.
Figure 10: Confidence score distribution using Llama-3-8B with the CoT method.
Figure 11: Confidence score distribution using Llama-3-8B with the MS method.
Figure 12: Confidence score distribution using Llama-3-8B with the SC-CA method.
Figure 13: Confidence score distribution in the RACE dataset.
Figure 14: Confidence score distribution in the Reddit Corpus (small) dataset.
Figure 15: Confidence score distribution in the WikiText dataset.
Figure 16: Confidence score distribution in the Reddit Corpus (small) dataset focusing on utterance of confidence.
Appendix N Calibration
Calibration is a key aspect of safety-critical systems, requiring that model confidence predictions are closely aligned with the level of predictive performance. Calibration error, which results from a mismatch between the accuracy of the model on sets of data at given levels of confidence, is quantified using a degree of metrics. For this analysis, we use the widely used binning-based metric, expected calibration error (ECE) [pakdaman2015ece], to measure the degree of calibration error induced post-attack. We centre this analysis on the results for GPT-4o and Llama-3-70B given their very high level of performance and hence larger capability of being miscalibrated through adversarial attacks leading to lowered confidence scores.
Table 18 compares how the ECE is affected post TB and SSR attacks across the main four tested CEMs. A value of 10 was used for the number of bins. We can see how the models have very high calibration both before (O.) and after (Adv.) undergoing attacks, and that in most cases for GPT-4o there is a increase in ECE on adversarial examples. In contrast, ECE is more mixed with Llama-3-70B with inconsistent patterns of miscalibration across different cases. This likely means that Llama is more overconfident and hence decreasing confidence through attacks can appear to reduce ECE. Nevertheless, in most cases the impact is significant between the original and adversarial scenarios, and thus warrants close attention. Figures 17 and 18 both show calibration diagrams for GPT-4o and Llama-3-70B respectively to reveal the general miscalibration patterns across all CEMs and attack types. Overall, as expected, the adversarial data has a high mismatch between accuracy and confidence in low confidence bins as a whole.
We also collect ECE results in Table 19 for the Likert Scale experiments from Appendix I. We observe similar results, although we do observe how ECE post attack increases here in most cases given that the Likert Scale based confidence is less highly confident overall, showing the harm to miscalibration when the model is not as overconfident.
Although the miscalibration is generally high prior to the confidence attacks, we stress that evaluating the robustness of these models’ confidence behaviour is crucial. Despite most CEMs producing verbal confidence that do not have ideal calibration, previous work has shown that LLM generated confidence scores do follow relative ordering by demonstrating that there is a statistically significant difference in LLMs (such as the tested GPT-4o and Llama-3-70B models here) verbal confidence level on correct answers versus incorrect answers [Omar2024.08.11.24311810] whereby correct answers contain a higher level of confidence on average. By attacking confidence, we can affect this ordering and decouple LLMs relative preference in confidence to true answers, which highlights issues in model honesty and perception of veracity. In sum, if we wish to create well-calibrated models through methods such as recalibration, we must also consider how vulnerable they are to adversarial attacks and address these vulnerabilities as well, otherwise we cannot ensure that models will be calibrated across all realistic inputs.
In addition, real-world production models and methods we detailed previously that are currently using verbal confidence in industry are vulnerable irrespective of the calibration. Furthermore, we emphasize that we also test CEMs designed to improve calibration, most notably MS, and our results show that while MS is far less overconfident and better calibrated in most scenarios, it is more vulnerable to attacks in terms of error rate and changes in confidence, hence better calibrated models are no less vulnerable than their miscalibrated versions. Lastly, we note that many important works such as in 10.5555/3305381.3305518 have shown that a large amount of vision models and LLMs are heavily miscalibrated and despite this many have studied adversarial attacks against these models, both traditional and confidence-based such as in obadinma2024calibration.
Table 18: ECE results for GPT-4o and Llama-3-70B.
Base CoT MS SC
O. ECE Adv. ECE O. ECE Adv. ECE O. ECE Adv. ECE O. ECE Adv. ECE
Llama-3-70B TQA VCA-TB 0.300 0.187 0.278 0.262 0.194 0.163 0.304 0.315
SSR 0.311 0.300 0.296 0.309 0.082 0.043 0.285 0.334
MQA VCA-TB 0.139 0.141 0.204 0.220 0.214 0.272 0.195 0.211
SSR 0.159 0.233 0.200 0.325 0.215 0.181 0.217 0.273
SQA VCA-TB 0.332 0.311 0.338 0.330 0.273 0.165 0.353 0.326
SSR 0.331 0.303 0.338 0.347 0.289 0.136 0.361 0.341
GPT-4o TQA VCA-TB 0.225 0.223 0.269 0.274 0.107 0.192 0.248 0.274
SSR 0.228 0.241 0.264 0.291 0.103 0.134 0.167 0.227
MQA VCA-TB 0.102 0.119 0.087 0.130 0.197 0.251 0.083 0.152
SSR 0.116 0.185 0.097 0.229 0.211 0.152 0.096 0.165
SQA VCA-TB 0.307 0.250 0.316 0.259 0.257 0.154 0.323 0.268
SSR 0.304 0.247 0.332 0.254 0.264 0.115 0.325 0.231
Table 19: ECE results for GPT-3.5 and Llama-3-8B on TQA when using Likert Scale based confidence.
Base CoT MS SC
O. ECE Adv. ECE O. ECE Adv. ECE O. ECE Adv. ECE O. ECE Adv. ECE
Llama-3-8B VCA-TF 0.47 0.46 0.51 0.50 0.22 0.30 0.37 0.41
VCA-TB 0.47 0.48 0.51 0.49 0.22 0.29 0.42 0.41
SSR 0.47 0.49 0.51 0.41 0.22 0.25 0.42 0.30
Ty. 0.47 0.44 0.51 0.47 0.22 0.22 0.39 0.36
GPT-3.5 VCA-TF 0.20 0.24 0.23 0.23 0.14 0.20 0.24 0.29
VCA-TB 0.19 0.23 0.22 0.24 0.12 0.13 0.23 0.31
SSR 0.19 0.26 0.23 0.27 0.14 0.16 0.21 0.25
Ty. 0.19 0.16 0.23 0.36 0.16 0.16 0.22 0.26
Figure 17: GPT-4 calibration diagrams across all attack types and CEMs. Left: Original Confidence Score Distribution Right: Adversarial Confidence Score distribution. Red bars are the average bin confidences, and blue bars are the accuracies within the bins.
Figure 18: Llama-3-70B calibration diagram across all attack types and CEMs. Left: Original Confidence Score Distribution Right: Adversarial Confidence Score distribution. Red bars are the average bin confidences, and blue bars are the accuracies within the bins.
Appendix O Brier Score Analysis
Table 20: Brier Score Results using the different types of attacks (against user queries) on SQA. Note in the majority of cases ( 80%) the confidence attacks increase the Brier score, showing miscalibration is induced. In most cases, the Brier score pre-attack is not severe (e.g., around 0.2) showing that verbal confidence is not heavily miscalibrated originally.
VCA-TF VCA-TB SSR Ty.
Pre Post Pre Post Pre Post Pre Post
Llama-3-8B SQA Base 0.51 0.52 0.51 0.51 0.51 0.42 0.51 0.42
CoT 0.07 0.15 0.07 0.19 0.07 0.20 0.07 0.28
MS 0.26 0.30 0.26 0.29 0.26 0.32 0.26 0.31
SC-CA 0.25 0.27 0.23 0.30 0.26 0.35 0.23 0.30
GPT-3.5 SQA Base 0.33 0.32 0.31 0.30 0.31 0.37 0.32 0.41
CoT 0.25 0.28 0.24 0.27 0.23 0.34 0.22 0.44
MS 0.23 0.24 0.23 0.28 0.22 0.30 0.23 0.32
SC-CA 0.23 0.30 0.22 0.29 0.22 0.37 0.23 0.41
To further analyze the calibration of verbal confidence we calculate Brier scores [VERIFICATIONOFFORECASTSEXPRESSEDINTERMSOFPROBABILITY] across different types of attacks and CEMs to determine whether CEMs possess a reasonable level of calibration according to this metric and whether attacks can further induce miscalibration. Brier score assesses both discriminatory accuracy and calibration and is a popular metric for calibration analysis. In Table 20 we present our results and show that Brier score in most cases pre-attack is reasonable and that in the vast majority of cases VCAs induce further miscalibration, showing the harm of these attacks.
Appendix P Confidence Score Change when Model Predicts Same Answer
Table 21: Δ affected confidence on adversarial attacks that do not flip the label.
Base CoT MS SC
Δ Aff. Cf. Δ Aff. Cf. Δ Aff. Cf. Δ Aff. Cf.
Llama-3-8B TQA VCA-TF 0.2 0.0 7.7 2.2
VCA-TB 0.9 0.0 9.4 3.1
SSR 3.6 2.1 12.1 3.6
Ty. 11.7 22.5 4.8 1.5
MQA VCA-TF 1.1 0.1 10.3 1.5
VCA-TB 1.1 0.3 12.8 2.0
SSR 3.6 0.2 16.8 2.8
Ty. 14.3 32.9 3.7 0.3
SQA VCA-TF 9.1 3.7 17.8 5.6
VCA-TB 7.6 3.8 21.5 14.5
SSR 16.3 4.4 17.4 12.0
Ty. 52.5 20.3 6.2 2.8
GPT-3.5 TQA VCA-TF 0.1 0.4 10.3 1.7
VCA-TB 0.1 0.5 12.8 1.6
SSR 0.4 1.6 16.8 2.9
Ty. 1.3 5.1 3.7 0.9
MQA VCA-TF 0.2 0.6 17.8 0.8
VCA-TB 0.2 0.9 21.5 0.9
SSR 1.3 1.9 17.4 2.3
Ty. 4.4 3.5 6.2 1.5
SQA VCA-TF 0.7 0.6 7.7 2.0
VCA-TB 0.8 1.1 9.4 3.6
SSR 1.2 3.9 12.1 5.3
Ty. 1.6 9.2 4.8 9.7
We analyze how confidence scores change post attack on queries in cases when the model predicts the same answer compared to the original (an unsuccessful conventional adversarial attack) to ascertain how much damage occurs in these cases, focusing on whether reduction to confidence scores is still great in cases where the overall adversarial attack is unsuccessful. In Table 21 we run this analysis over all of the main datasets and attacks and CEMs for our Llama-3-8B and GPT-3.5, focusing on attacks on user questions. The results reveal that there is large variance between the Δ Aff. Cf. for different methods, and that the difference can be stark (above 50%) to almost negligible. In most cases, the Δ Aff. Cf. for these examples where the answers were unchanged is smaller than the overall Δ Aff. Cf. when considering all samples seen in Table 2 such as in the case of TruthfulQA with Llama using the Base method (0.2 vs. 19.3), meaning that the examples where the label is flipped are more susceptible to significant confidence reduction.
Appendix Q Attacks Optimized for Pure Confidence Degradation
Given that our adversarial attacks are based on attacking verbal confidence and trying to induce answer changes, in this section we test the performance of pure confidence attacks i.e., when the focus is on manipulating confidence scores without changing the original predicted answers. This can allow us to better determine where the weakness of confidence elicitation lies insofar as we can see whether it is verbal confidence that is specifically vulnerable or whether there is a general vulnerability that leads to different model outputs.
As jailbreak-based attacks cannot guarantee that original answers are preserved on downstream examples, we choose the most effective perturbation-based attack method, SSR and change the attack objectives so that effective perturbations cannot also change the original generated answer, and observe how the difference in average confidence change is affected. Table 22 shows the results of these experiments. We observe how in some cases it appears allowing the attack algorithm to change the answer leads to larger drops in confidence, while in other cases the average confidence level on adversarial examples is higher than before. SC-CA in particular seems to fare poorer when not allowed to induce label flips with the majority of cases the confidence drop is lower than the original attack algorithm. In all cases though, the difference is minor between both variants of the attack algorithm, showing that confidence vulnerability is not strongly tied to whether the model’s answer is changed as a result of adversarial perturbations if allowed to optimize further beyond achieving a label flip. In essence, given the previous section this means that sensitive examples where label flips tend to occur also remain very vulnerable to continued confidence attacks even if a label flip does not occur.
Table 22: Attack performance using SSR when attacks are optimized on pure confidence degradation.
Cf. Adv. Cf. %Aff. #Iters Δ Aff. Cf.
Llama-3-8B TQA Base 87.1 82.6 17.5 7.1 25.6
CoT 86.7 83.9 6.5 6.8 42.6
MS 61.0 49.5 58.0 7.3 19.8
SC-CA 87.2 79.9 46.5 6.6 15.8
MQA Base 87.7 83.8 15.0 9.2 25.8
CoT 89.1 86.3 14.0 13.5 20.0
MS 58.1 42.6 61.5 11.0 25.2
SC-CA 87.7 79.0 54.5 9.2 16.0
SQA Base 43.2 24.7 32.0 6.6 57.8
CoT 92.4 86.4 23.5 5.9 25.5
MS 83.6 68.0 70.0 6.8 22.3
SC-CA 89.3 76.2 62.5 5.6 21.0
Avg. - 9.2 38.5 8.0 26.4
GPT-3.5 TQA Base 94.3 93.8 7.5 11.4 7.3
CoT 92.2 90.8 18.0 7.6 7.5
MS 69.6 58.7 64.0 7.5 17.0
SC-CA 93.2 89.9 61.0 6.9 5.4
MQA Base 94.4 93.6 7.5 11.3 10.7
CoT 95.7 94.2 26.5 9.2 5.8
MS 64.0 52.4 63.5 9.6 18.4
SC-CA 96.3 93.8 48.5 9.5 5.2
SQA Base 97.0 95.7 24.5 6.4 5.2
CoT 95.5 93.6 23.0 6.4 7.9
MS 84.0 65.6 71.5 6.7 25.7
SC-CA 96.0 90.1 59.0 6.3 10.1
Avg. - 5.0 39.5 8.3 10.5
Appendix R Detailed Defence Results Against ConfidenceTriggers
Table 23: Perplexity filter and LLM-Guard results against ConfidenceTriggers-AutoDAN on Llama-3-8B. A negative Δ shows average confidence increased post filter.
SQA TQA MQA
% fil. Δ Cf. %MI % fil. Δ Cf. %MI % fil. Δ Cf. %MI
Perp. Base 29.6 -0.6 0.0 0.0 0.0 0.0 7.7 0.1 0.0
CoT 21.1 1.1 0.0 21.4 -3.7 0.0 21.1 2.1 0.0
MS 49.0 -0.4 0.0 7.8 0.2 0.0 74.5 -0.3 0.0
Guard Base 37.7 -3.0 2.0 37.5 -2.6 2.2 20.5 -0.5 12.4
CoT 0.0 0.0 0.6 100.0 28.8 2.3 0.0 0.0 0.0
MS 0.0 0.0 0.0 0.0 0.0 0.9 0.0 0.0 0.0
We test methods from the three main categories of jailbreak-based methods: Self-Processing Defences, Additional Helper Defences, and Input Permutation Defences [xu-etal-2024-comprehensive]. Firstly, we test the popular the self-processing perplexity filter defence approach from jain2023baselinedefensesadversarialattacks. Here, we optimize a perplexity threshold for each dataset/CEM combination that results in no normal samples being filtered but can filter any text with abnormal perplexity (such as those containing trigger prompts). We also test an additional helper defence using the LLM-Guard toolkit [goyal2024llmguardguardingunsafellm], using a detection based approach to filter out prompts with a high level of gibberish text. Finally, we include two input permutation defences, the Paraphrase defence [jain2023baselinedefensesadversarialattacks], and a modified version of SmoothLLM [robey2023smoothllm]. The former uses an additional LLM (GPT-3.5) to rephrase an input prompt containing a trigger phrase. The latter samples multiple variants of an input prompt and performs permutations (character swapping) before feeding the modified prompts to the LLM and using majority voting to decide on the answer and averaging the confidence across the different prompts to attain the final confidence.
We conduct our experiments on the same set of data as Table 2. We test mitigation strategies using deterministic CEMs (Base, CoT, and MS) since they are single sample and the minimal randomness allows us to determine clearly what strategies are most effective. Table 23 presents the results of the perplexity filter defence. We show the percentage of user queries that get filtered (%fil.) (only including those that the trigger originally led to a confidence decrease) and the resulting loss in average confidence on those samples (Δ Cf.) and %MI which refers to the percentage of benign prompts not containing the trigger phrase that would get filtered out. We can see that in most cases the majority of these samples do not get filtered and the samples that do not lead to a big drop in average confidence (so mainly queries with low confidence drops get filtered). Similarly, Table 23 presents the results for LLM-Guard. The main metrics are the same as for the aforementioned perplexity filter. Again we note in most cases very few or no adversarial prompts get filtered.
Table 5 presents the results for the two input permutation defences. We show the difference in average confidence provided by the LLM between the original set of data and those with the permuted prompts without any trigger phrase (Δ Cf. O.) and with it (Δ Cf. A.). Additionally, we note what percentage of permuted examples with originally benign (% Aff.) and trigger-included prompts (% Adv. Aff.) had an altered generated confidence level compared to the original un-permuted version. Finally, we note the average decrease in confidence after input permutation on the trigger-included prompts (Δ Adv.). Our primary findings are that the permutation-based defences lead to significant changes in confidence both when the original prompt does and does not include a trigger phrase (with differences of up to  43%). The majority of examples in most cases have their confidence altered as a result of these defences and there are very few cases where minimal changes of confidence occur after permutation of an adversarial prompt signalling that these defences are largely ineffective.
Nevertheless, without even considering defences such as perplexity-based ones that are popular for trigger-based attacks, ConfidenceTriggers can still present issues in specific applications. For example, if an adversary gains access to a custom LLM and supply a compromised system prompt (or demonstration) with ConfidenceTriggers-AutoDAN that is difficult to notice, then an external user using this model with their own queries would be affected. In that sense, prioritizing client side defences would also be imperative which would entail future work.
GPT-3.5
We collect additional defence results against ConfidenceTriggers-AutoDAN using GPT-3.5. We test three main defences (since the perplexity filter is not feasible without access to logits): Paraphrase, SmoothLLM, and GPT-4-Filtering (an original approach similar in idea to the LLMGuard defence). We present the results for the first two in Table 24 using the same approach and metrics as we used for the Llama-3-8B results. As we have observed before, the input perturbation based defences lead to major changes in average confidence both when perturbing the original system prompts and those containing the triggers, meaning they are largely ineffective in minimizing the impacts on confidence. In fact, we observe up to 31% decrease in average confidence when applying the paraphrase defence on prompts including the trigger phrases. The results for GPT-4-Filtering are in Table 25 with the same metrics used for LLM-Guard. We note that only a small percentage of examples where the system prompts with trigger phrases led to a confidence decrease get filtered. Furthermore, we observe often high rates (up to  19%) of examples without any trigger phrase being filtered by this approach, showing that it is largely ineffective.
Base ConfidenceTriggers
Finally, we collect results using the perplexity filter on prompts with the base ConfidenceTriggers algorithm triggers. Table 26 demonstrates that such triggers, as established in previous works are vulnerable to perplexity filter defences given the unnaturalness of the random sequence of tokens. In this sense there is a trade-off between detectability and performance of the attacks.
Table 24: ConfidenceTriggers-AutoDAN Paraphrase and SmoothLLM defense results on GPT-3.5. Negative Δ’s show that average confidence increased post permutation.
SQA TQA MQA
Paraphrase
Δ Cf. O. Δ Cf. A. % Aff. % Adv. Aff. Δ Adv. Δ Cf. O. Δ Cf. A. % Aff. % Adv. Aff. Δ Adv. Δ Cf. O. Δ Cf. A. % Aff. %Adv. Aff. Δ Adv.
Base 5.8 7.9 64.0 84.7 8.7 -1.6 6.9 80.0 75.0 10.8 17.1 12.1 71.5 82.5 8.8
CoT 1.9 1.8 46.0 30.6 12.0 9.9 5.7 62.5 37.3 27.3 13.1 15.8 55.5 55.1 27.5
MS 6.6 11.3 81.0 65.8 17.2 17.3 20.9 90.5 57.9 27.3 20.8 31.2 95.0 60.0 41.8
SmoothLLM
Base 3.1 6.0 62.0 90.0 5.5 4.3 6.3 63.0 75.5 5.4 4.9 8.1 72.0 91.0 6.3
CoT 2.1 2.7 50.5 56.0 4.7 3.4 4.8 63.5 71.0 5.0 2.7 3.6 66.0 71.5 3.0
MS 4.8 14.7 96.0 88.5 11.1 5.7 12.4 99.0 79.5 8.6 5.6 8.5 98.5 75.5 4.8
Table 25: LLM-Guard Results for GPT-3.5. %MI refers to percentage of non-adversarial samples classified as adversarial. Negative Δ’s show that average confidence increased post filter.
SQA TQA MQA
% filtered Δ Cf. %MI % filtered Δ Cf. %MI % filtered Δ Cf. %MI
Base 2.0 0.0 3.9 7.1 -0.2 2.7 24.1 -0.1 0.0
CoT 0.0 0.0 1.4 2.5 -0.1 2.5 2.2 -0.1 0.9
MS 14.2 -0.5 11.1 6.4 0.1 18.7 4.5 0.0 1.1
Table 26: Perplexity filter results using base ConfidenceTriggers using Llama-3-8B model.
SQA TQA MQA
% filtered Δ Cf. % filtered Δ Cf. % filtered Δ Cf.
Base 100.0 0.0 99.7 0.0 95.7 -0.05
CoT 100.0 0.0 86.3 -3.7 100.0 0
MS 100.0 0.0 92.5 0.2 88.5 -0.01
Appendix S Jailbreak Defence Descriptions
Perplexity Filter [jain2023baselinedefensesadversarialattacks]
We use a naive filter based on the text perplexity of the system prompt. The text perplexity of sequence of tokens is defined as the averaged negative log likelihood of the tokens. The log perplexity (pplppl) is thus defined for a whole sequence of tokens $\mathbf{\mathcal{X}}=\{x_{1},\dots,x_{N}\}$ by $log(ppl)=\frac{-1}{|\mathbf{X}|}\sum_{i}P_{i}\;\mbox{log p}(x_{i}|x_{0:i-1})$. Given that the log likelihood of a token in a sequence represents its probability of appearing in that context, a high negative log likelihood would mean that the tokens in a sequence are likely unnatural or manipulated. For our implementation, we set a threshold such that any full prompt with a perplexity higher than the threshold is considered filtered. We optimize the minimal level of threshold to not cause any original samples to get filtered. Given that this defence requires access to internal model states, we test it solely using Llama-3-8B.
LLM-Guard [goyal2024llmguardguardingunsafellm]
is a popular toolkit containing many types of scanners (trained models) that attempt to identify malicious text like prompt injections, toxic language, and confidential information. We utilize the gibberish scanner that can identify and filter out gibberish or nonsensical inputs such as those containing trigger phrases. In essence we feed the scanner with the full prompt (system prompt, one-shot example and user query included) and see whether it flags the prompt. We use the default threshold of 0.9 and any text with a gibberish score above is filtered. We do not use the prompt injection scanner as we find that fails to identify any of our prompts containing triggers.
Paraphrase Defence
We follow the approach of jain2023baselinedefensesadversarialattacks which follows the protocol of kirchenbauer2024on for this defence. We instruct GPT-3.5 to paraphrase the system prompt according to the instruction “Please paraphrase the following sentences in under 200 words and only output the revised version, the sentence is:”. We use a temperature of 0.7 for the paraphrased output and limit the number of tokens to 400. The goal is that by paraphrasing the impact of a system prompt containing a trigger is minimized.
SmoothLLM-Confidence-Attack [robey2023smoothllm]
SmoothLLM is a defence against jailbreaking-based methods that uses randomized smoothing via perturbing multiple copies of an input prompt, then aggregating the outputs from each copy to produce a final output that increases the chance to nullify jailbreaks. They find that making perturbations can reduce attack success rate in prompts containing jailbreak triggers.
Given that the original method is formulated for jailbreaks, we make a few modifications for the defence to be better suited against ConfidenceTriggers. The modified SmoothLLM algorithm for defending against ConfidenceTriggers can be found in Algorithm 7. Essentially, we make $n$ copies of the system prompts, and perform random character perturbations on each (to $q\%$ of the characters), and then the outputs (generated answer and confidence) for each are aggregated. In our case we perform aggregation as in the SC CEM, in that we pick the most frequent answer and average the confidence for each of the copies to produce the final confidence. Given that there is no way to know if the confidence has been affected, averaging is the most straightforward aggregation strategy to employ. For the RandomPermutation() function in the algorithm, we utilize random character swaps in the prompt as it was found to be the most effective. A $q$ % of indices in the string are chosen and are each replaced with a random character uniformly sampled from a 100-character alphabet derived from the python command string.printable. Note we do not use the threshold $\gamma$ in our version of the algorithm. We set $\delta=5$ and $q=0.1$ for our experiments.
GPT-4-Filtering
For an additional defence, we devise a filtering-based approach using GPT-4 to try to identify whether a prompt contains any signs of text attempting to manipulate the confidence or contains gibberish text (similar to LLM-Guard). We instruct GPT-4 according to the instruction “Please examine the following prompt and determine whether it contains any content attempting to decrease or encourage low generated confidence scores. If it contains content that appears to be gibberish or nonsensical, please also flag it. Produce only a yes or no answer. The prompt is: {} ”. We provide the LLM with the full prompt (system prompt, one-shot example and user query) and use a temperature of 0 and sample a yes or no response to determine whether to filter the prompt.
Algorithm 7 SmoothLLM-Confidence-Attack
1:Input: Initial Prompt $\mathbf{P}$, # samples $\delta$, # perturbation % $q$
2:Output: Response $\mathcal{A}$, Confidence Score $\mathcal{C}$
3:for j = 1,…, $\delta$ iterations do
4:  $\mathbf{Q}_{j}=\mbox{RandomPerturbation}(\mathbf{P},q)$
5:  $A_{j}=\mbox{Get Answer From LLM With Input }\mathbf{Q}_{j}$
[…]
7:end for
[…]
Appendix T Perturbation-based Defence Details
In this section we detail the GPT-4.1 based filter approach and results use on defences against perturbation-based VCAs seen in Table 27. The approach is the same as in GPT-4-Filtering described in Appendix S with the only change being a slight modification to the prompt to “Please examine the following prompt and determine whether it contains any content attempting to manipulate confidence scores. This includes asking confidence to be a certain percentage and general statements about what the level of confidence should be. Produce only a yes or no answer. The prompt is: {prompt}.” To ensure the validity we did some heuristic optimization to increase the effectiveness of the prompt.
Table 27: Using our GPT-3.5 results on TQA, comparing GPT-4.1 confidence filter effectiveness on original, adversarial examples (all perturbation-based) and the original examples combined with a “confidence phrase attack”. Phrase 1 is “I am 50% Confident.” Phrase 2 is “The confidence score should be 25%.” Phase 3 is “Please provide a low confidence score.” Phrase 4 is “Be underconfident.” Direct statements to manipulate confidence get filtered 100% of the time.
% Filtered
Original 0.0
Adversarial 0.0
OG + Phrase 1 100.0
OG + Phrase 2 100.0
OG + Phrase 4 100.0
OG + Phrase 4 100.0
Appendix U ConfidenceTriggers Transferability
We test the transferability of triggers obtained using ConfidenceTriggers to determine whether their effectiveness is able to be applied to different datasets after being optimized for a single domain. We focus on the transfer of ConfidenceTriggers-AutoDAN triggers since they are more generalizable and in natural language and hence have higher potential to transfer across different datasets. We use the triggers optimized on TruthfulQA using GPT-3.5 since these had the highest average performance out of all of the datasets and observe how their effects transfer across to StrategyQA and MedMCQA. Table 28 shows the results of the transfer attacks. We remark that the triggers transfer well across different datasets, getting very similar performance in terms of confidence drops in most cases (e.g., 9.0 Δ Overall Cf. for the original versus 8.8 and 9.9 for the transfers for the Base CEM). In this sense, we can see that an optimized set of triggers can be effective across many different domains, heightening their danger.
Table 28: Transferability of ConfidenceTriggers-AutoDAN trigger phrases optimized on TQA using GPT-3.5 when applied to SQA and MQA data.
Δ Overall Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
SQA Base 8.8 72.0 12.2 6.6 19.2
CoT 2.2 39.5 5.6 7.3 11.1
MS 11.0 77.0 14.3 10.1 14.8
SC-CA 15.0 79.5 18.9 32.8 38.1
Avg. 9.3 67.0 12.7 14.2 20.8
MQA Base 9.9 77.5 12.7 20.0 21.1
CoT 2.9 48.0 6.1 7.3 15.6
MS 6.9 68.5 10.0 3.9 19.4
SC-CA 2.7 69.0 3.9 10.3 24.3
Avg. 5.6 65.8 8.2 10.4 20.1
Appendix V AUROC Analysis
We provide AUROC scores to give a better idea of whether perturbed confidence scores still correlate with answer correctness after undergoing perturbation-based VCAs. In Tables 29-32, we show the results of this based on our main results on SQA in Table 2, Table 9 (and Figure 2), Table 3/10, and Table 4 respectively. We focus on SQA since the predictions are binary and we are easily able to obtain the probabilities for the non-predicted label. Due to the lack of predicted probabilities for non-predicted classes in this multiple choice question answering setup, attaining a valid AUROC score for the other datasets requires extensive compromised (e.g., using a one-versus-all setup) that make interpreting the score difficult. We find that in almost all cases AUROC scores lean more towards random guessing (a score of 0.5) post VCAs, showing that VCAs are effective at harming both model confidence and predictive performance.
Table 29: AUROC scores on SQA based on results featured in Table 2.
VCA-TF VCA-TB SSR Ty.
Pre Post Pre Post Pre Post Pre Post
Llama-3-8B Base 0.40 0.42 0.40 0.44 0.40 0.56 0.40 0.56
CoT 0.57 0.58 0.57 0.54 0.57 0.54 0.57 0.52
MS 0.67 0.57 0.67 0.59 0.67 0.51 0.67 0.57
SC 0.70 0.64 0.71 0.62 0.69 0.54 0.75 0.64
GPT-3.5 Base 0.61 0.62 0.63 0.64 0.65 0.56 0.63 0.48
CoT 0.70 0.67 0.70 0.69 0.72 0.62 0.73 0.45
MS 0.70 0.67 0.69 0.61 0.73 0.56 0.70 0.52
SC 0.74 0.66 0.76 0.67 0.75 0.56 0.72 0.44
Table 30: AUROC scores on SQA based on results featured in Figure 2/ Table 9.
VCA-TF VCA-TB SSR Ty.
Pre Post Pre Post Pre Post Pre Post
Demo. Llama-3-8B Base 0.40 0.47 0.40 0.46 0.40 0.56 0.40 0.39
CoT 0.57 0.61 0.57 0.62 0.57 0.57 0.57 0.59
GPT-3.5 Base 0.61 0.64 0.65 0.63 0.65 0.61 0.62 0.63
CoT 0.72 0.72 0.72 0.74 0.73 0.68 0.70 0.70
Sys. Llama-3 8B Base 0.40 0.45 0.40 0.43 0.40 0.54 0.40 0.58
CoT 0.57 0.49 0.57 0.54 0.57 0.44 0.57 0.56
GPT-3.5 Base 0.61 0.62 0.62 0.65 0.64 0.65 0.60 0.67
CoT 0.72 0.76 0.72 0.76 0.72 0.73 0.72 0.72
Table 31: AUROC scores on SQA based on results featured in Table 3/10.
VCA-TB SSR
Pre Post Pre Post
Llama-3-70B Base 0.77 0.67 0.77 0.57
CoT 0.82 0.73 0.82 0.56
MS 0.85 0.74 0.84 0.56
SC 0.82 0.75 0.81 0.66
GPT-4o Base 0.85 0.73 0.86 0.53
CoT 0.87 0.79 0.89 0.54
MS 0.87 0.71 0.87 0.53
SC 0.87 0.72 0.88 0.58
Table 32: AUROC scores on SQA based on results featured in Table 4.
Random Tokens ConfidenceTriggers ConfidenceTriggers-AutoDAN
Pre Post Pre Post Pre Post
Llama-3-8B Base 0.45 0.54 0.51 0.46 0.47 0.52
CoT 0.54 0.51 0.52 0.56 0.56 0.52
MS 0.45 0.50 0.48 0.54 0.57 0.53
SC 0.53 0.59 0.52 0.47 0.56 0.54
GPT-3.5 Base 0.64 0.62 0.67 0.58 0.67 0.65
CoT 0.76 0.71 0.73 0.67 0.72 0.67
MS 0.74 0.74 0.73 0.73 0.75 0.73
SC 0.75 0.75 0.76 0.72 0.72 0.69
Appendix W Prompt Examples
Figure 19 portrays an examples of a prompt that we use for the CEMs (in this case CoT). An example for Self-Probing can be seen in Figure 20. Additionally, we provide example one-shot demonstrations we utilize for MedMCQA (based on an example from the training set) that help steer the model to produce outputs in the correct format and allow it to effectively conduct the task in Figure 21.
Figure 19: An example prompt for CoT method from TruthfulQA dataset.
Figure 20: An example system prompt for Self-Probing method .
Figure 21: One-shot demonstrations examples for MedMCQA across different CEMs.
Appendix X ConfidenceTriggers Examples
Figure 22 gives an example of the kinds of trigger prompts both variants of ConfidenceTriggers generates.
Figure 22: Examples of trigger prompts generated using ConfidenceTriggers.